### v1

# Application of AI-based Models for Online Fraud Detection and Analysis

Abstract
[…]
Background Fraud is a prevalent offence that extends beyond financial loss, impacting victims emotionally, psychologically, and physically. The advancements in online communication technologies, allow for online fraud to thrive in this vast network, with fraudsters increasingly using these channels for deception. With the progression of technologies like Generative Artificial Intelligence (GenAI), there is a growing concern that fraud will increase and scale, using these advanced methods such as deep-fakes in phishing campaigns. However, the application of AI in detecting and analyzing online fraud remains understudied. This review addresses this gap by investigating AI’s role in analyzing online fraud using text data.
[…]
Methods We conducted a Systematic Literature Review (SLR) on AI and Natural Language Processing (NLP) techniques for online fraud detection. The review adhered to the PRISMA-ScR protocol, with eligibility criteria including language, publication type, relevance to online fraud, use of text data, and AI methodologies. Out of $2,457$ academic records screened, $350$ met our eligibility criteria, and $223$ were analyzed and included herein.
[…]
Results We report the state-of-the-art NLP techniques used to analyse various online fraud categories; the data sources used for training the NLP models; the NLP algorithms and models built; and the performance metrics employed for model evaluation. We find that the current state of research on online fraud is broken into the various scam activities that take place, and more specifically, we identify $16$ different frauds that researchers focus on. Finally, we present the latest and best performing AI methods used towards detecting online scam and fraud activities.
[…]
Conclusions This SLR enhances the academic understanding of AI-based detection methods for online fraud and offers insights for policymakers, law enforcement, and businesses on safeguarding against such activities. We conclude that focusing on specific scams lacks generalization, as multiple models are required for different fraud types. Furthermore, the evolving nature of scams limits the effectiveness of models trained on outdated data. We also identify issues in data limitation or training bias reporting from researchers, along with issues on model performance reporting, with some studies selectively presenting metrics, leading to potential biases in model evaluation.
Artificial Intelligence,
Natural Language Processing,
Online Fraud,
Systematic Literature Review,
keywords:
[…]
Systematic Review*
*NOTE: This manuscript is currently under peer review. Please be aware that the content may be subject to revision in future versions.
{abstractbox}
Background
Online fraud has emerged as one of the most pervasive and challenging threats in the digital age affecting individuals of all ages, businesses of different sizes, and governments. Defined broadly, online fraud is an umbrella term that involves acts of deception or deliberate impersonation, taking place on the Internet, towards the personal gain of the fraudster, often resulting in a financial loss for the victim ukfinance2023 . Other than financial losses, fraud can have a wide range of impacts on victims. These include emotional and psychological effects such as anger, fear, shame, depression, decrease in confidence, and trauma; impacts on physical and mental well-being; harms relationships and leads to loneliness and isolation; and causes negative changes in behaviour parliament_post_2020 . Although there is evidence suggesting that certain socio-demographic groups face higher risks of fraud–such as women aged 25-34 or 35-44 experiencing fraud at rates of $7.6\%$, and individuals in the highest income bracket at $8.3\%$–fraud is impacting individuals across all demographics parliament_post_2020 and sometimes in different ways. For example, victims earning £20,000 or less, those aged 65 and over, and female victims, reported impacts on their self-confidence, at rates of $46\%$, $43\%$, and $41\%$, respectively, compared to the overall victim rate of $35\%$.
For the year ending March 2023, the Crime Survey for England and Wales estimated 3.5 million fraud offences including online fraud  ons2023 . For example, advance fee fraud increased significantly highly from 60,000 to 391,000 offences compared to the year ending March 2020. This increase is largely due to society’s growing reliance on the Internet and digital platforms for everyday services, transactions and communications. According to The Office of Communication (Ofcom) ofcom2023 , $92\%$ of adults in the UK use the Internet for a wide variety of activities including communication, education, and entertainment. Activities such as banking, shopping, and socialising are increasingly happening via online platforms, expanding the landscape for fraudsters to exploit vulnerabilities or use these platforms to deceive victims. In 2020, online shopping scams made up $38\%$ of all reported scams worldwide, an increase of $6\%$ compared to pre-Covid-19 outbreak statista-ecommerce-fraud .
Online fraud encompasses a wide range of deceptive activities, including identity theft, phishing, advance fee fraud, romance scams, fraudulent investment scams, and more. It is important to highlight that there is no universally accepted definition of “online fraud,” and the term is often used interchangeably with the term “scam.” Legally, “fraud is defined as false representation to cause loss to another or to expose another to a risk of loss” ukfraudlegislation2006 , and scam is the process where criminals gain the trust of victims to deceive or cheat victims ukscamglossary2006 , through false representation and other means, so that the victim trust them, which in turn results in various kinds of losses for the victim. The National Fraud Authority published a literature review fraudtypologiesuk where they compared the distinction of the term fraud as defined by the amended Fraud Act 2006 ukfraudlegislation2006 and the typology produced by Levi levi2008definingfraud . They found fraud embraces a broad scope of different crimes, whereas scams often are focused on fraud against individuals and small firms. For example, different scams like advance fee, romance, tech support, etc., all fall under the fraud umbrella, but they are also deception methods, which are in part scams. Hence, in this SLR, we use both terms as the various scams are the different deception methods scammers use to trick victims, and fraud is the term that includes all scams.
Online frauds exploit the virtual nature of the Internet and the anonymity it provides to reach out to victims. This virtual environment, coupled with jurisdictional challenges (where offenders and victims may be in different regions of the world), makes fraud difficult to detect and prevent using traditional policing techniques. The complexity of online fraud is further heightened by its evolving nature, as fraudsters continuously adapt their techniques to bypass new security measures and exploit emerging technologies to target on new victims skidmore24 .
Given the scale and impact of online fraud, there is a need for new methods to detect and prevent such activities. The use of NLP, in combination with other AI, has been proposed for identifying, characterising, and detecting fraudulent patterns in applications like phishing 9795286 , fake job advertisement amaar2022detection and analysing scam patterns lwin2023supporting which could help develop preventive measures and mitigate risks of online fraud. However, understanding the current state of AI techniques in combating online fraud, the data sources used, the evaluation methods for AI models, and the specific types of fraud that are most prevalent remain a significant challenge. This is due to the constant emergence of new fraud activities that use various communication mediums and social engineering attacks, in an attempt by fraudsters to remain undetected. Therefore, there is a pressing need to shift from detecting and analysing the effects of fraud to early detecting emerging fraudulent activities online and new methods of social engineering.
This systematic literature review (SLR) aims to address these challenges by providing a comprehensive overview of the state-of-the-art AI techniques used to detect fraudulent activities online. More specifically, we explore the data sources widely used by researchers to study online fraud, how researchers evaluate the methodologies employed to assess AI models, and identify the most popular types of online fraud targeted by researchers. By synthesising findings from academic papers, this review aims to provide a thorough understanding of the current landscape of online fraud detection and prevention, highlighting gaps in existing research, and proposing directions for future studies.
Research Questions. The research question for the Systematic Literaure Review (SLR) is as follows:
  • RQ1: What is the state-of-the-art of AI techniques used to detect online fraud?
  • RQ2: What are the data sources that researchers use to analyse online fraud?
  • RQ3: How do researchers evaluate their AI models?
  • RQ4: What are the most popular fraud activities that researchers study that fraudsters use to approach their potential victims?
Although a wide number of studies have explored the application of AI for fraud detection and other types of cybercrime, we are not aware of any systematic literature reviews that have examined the application of AI models using text data. The focus of this SLR is AI-based models that study textual data to detect and gain insights related to online fraud. Thus, the focus of this study is to identify Natural Language Processing models that are used to detect online fraud.
1 Online Fraud and AI
We now discuss online fraud and various AI methodologies in detail for further reader information. Online fraud refers to any deliberate act of deception conducted over the Internet to obtain an unlawful or unfair gain. It involves exploiting online platforms, services, and technologies to deceive individuals or organisations for financial, personal, or material gain. Online fraud can take many forms, each characterised by the method of deception and the medium used to perpetrate the fraud.
1.1 Fraud Categories
The list of online fraud activities is extensive and constantly evolving with new types or sub-types emerging. Developing a comprehensive taxonomy or classification for all online fraud activities requires special attention, which is beyond scope of this work. Below we outline some of the well-known online fraud types.
  • Phishing is the process where fraudsters impersonate representatives of legitimate organisations or acquaintances of the targeted victim to trick them into providing personal information such as usernames, passwords, credit card details, or bank account details. This activity can be done through various mediums, like email, phone calls (aka Vishing), SMS (aka Smishing), and any other way of online communication. There are various phishing scams that surfaced over the years, including the Royal Mail scam royalmailscams , banking scams barcleysscams , HMRC scams hmrcscams , and many others. Notably, phishing scams often include deceptive web addresses created by cyber-criminals to trick victims into believing they are visiting legitimate websites. The primary goal of these URLs is to steal personal data including usernames, passwords and credit card details for financial gain.
  • Fake Reviews are deceptive or fraudulent reviews created to mislead potential customers about the quality, reliability, or legitimacy of a product, service, or app. On fraudulent e-commerce websites and app stores, fake reviews play a crucial role in tricking victims into trusting and using fraudulent apps or purchasing substandard or non-existent products. This leads to potential victims trusting fraudulent websites, services, or apps, providing them with their credit card details for a purchase, that leads the victim to a vulnerable position paul2021fake .
  • Recruitment Fraud is a type of online scam where fraudsters pose as legitimate employers or recruiters to deceive job seekers. The primary goal of these scams is to receive “fees” for a job application, steal personal information, extort money, or exploit the victim in some other way. This type of fraud preys on individuals seeking employment, often targeting those who are most vulnerable or desperate for work mehboob2021smart .
  • Romance fraud (aka romance scams or dating scams) involves fraudsters creating fake profiles on dating websites, social media, or other online platforms to deceive victims into believing they are in a genuine romantic relationship. The primary objective is to exploit the victim’s emotions to extort money, personal information, or other benefits. This is an elaborate scam that is extremely difficult to detect, since it is also underreported due to victims feeling ashamed and hurt for being a victim from someone they considered to be a romantic partner coluccia2020online . In this scams, fraudsters communicate with victims for a long time before presenting them with an “investment opportunity” or requesting their financial aid. Romance scams are closely related to the Cryptocurrency Pig Butchering scams cross2023romance , where victims are gradually lured into making increasing contributions over a long period of time, usually in cryptocurrency, to a fraudulent scheme ordekian2024sinister .
  • Fraudulent Investment include scams where fraudsters promise victims significant winnings or lucrative opportunities vasek2015there . These scams are usually associated with the romance scams discussed above. Once the victims try to withdraw their “winnings” the scammers will extort them by asking for “fees” and “taxes” to be paid in advance. The promised benefits and winnings never materialize, and the initial investment sums and fees are lost agarwal2023defi . Fraudulent investment is the umbrella that covers Cryptocurrency Pig Butchering scams explained above, and various Ponzi schemes vasek2019analyzing where early investors greatly benefit from the investments of later investors, also known as pyramid schemes.
  • Crypto market manipulation involves artificially increasing or decreasing the price of cryptocurrencies to achieve financial gain. It often involves coordinated efforts by individuals or groups to manipulate the market to create false perceptions of supply, demand, or market sentiment. Some common techniques used in crypto market manipulation include: Pump and Dump which inflates the price of a cryptocurrency through misleading or false statements (pumping), encouraging others to buy, and then selling off the cryptocurrency at a profit once the price has been pumped up (dumping); Wash Trading occurs when a trader buys and sells the same cryptocurrency simultaneously to create deceptive activity on the market; Spoofing involves placing significant buy or sell orders to withdrawn them before execution to mislead perceptions related to the market demand or supply;Front-running involves placing orders ahead of a large trade that is known to occur, to benefit from the subsequent price movement caused by the large trade; and many others hamrick2021examination .
  • Fraudulent e-commerce involves deceptive practices or scams conducted through online e-commerce platforms. These scams aim to exploit the digital payment systems to deceive consumers or businesses via paying for a fraudulent product or service.
  • Fraudulent crowdfunding refers to the misuse of crowdfunding platforms to deceive donors or backers, often by providing false or misleading information about the nature, purpose, or outcome of a crowdfunding campaign. Crowdfunding itself is a method of raising money from a large number of people via online platforms to fund projects, products, or causes cumming2021disentangling . A fraud similar to crowdfunding is Charity Fraud and Disaster Scams where scammers seek donations for organisations that do not exist or do no or little work at all. These scams are particularly common after high profile disasters as criminals often use tragedies to exploit people who are looking to donate charityfraud .
  • Gambling Fraud is any illegal activity that is intended to cheat players or an online gambling platform. Fraudsters manage to trick victims and platforms in different ways, including rigged games, fake websites (phishing URLs described above), account takeovers (via stealing legitimate users’ access codes), and via creating fake apps with fake reviews, as discussed above to gain the trust of users. Online gambling fraud can happen on multiple platforms and involve wide variety of games inclduing casino scams, sports betting scams, and lottery scams  hong2022analyzing .
  • Tax Scams occurs when scammers falsify information regarding pending tax money, or via maliciously impersonating tax officials to trick individuals or business entities to willfully pay them ‘fees” brody2014income . Scams similar to tax scams are Council Tax Scams, various Utility bill scams, Insurance Scams, etc., where scammers impersonate officials and trick the victim that they own money. These scams fall under the umbrella of Phishing as they often take place via SMS, phone calls, or emails.
  • Pension Scams are similar to Tax Scams. Scammers aim to make money through fees, direct access to pension savings or by receiving investments pensionscams .
Overall, in this section we briefly described some well known scams and online fraud activities. The complexity and interconnected nature of scams and frauds make it challenging to categorise them under a single typology. Phishing scams, for instance, serve as a broad umbrella that currently encompasses phishing conducting using various methods like vishing (voice) and smishing (sms), yet they can also be integral parts of investment scams when scammers develop phishing websites to gain the trust of victims. Similarly, scams often involve tricking victims through deceptive tactics, ultimately leading to defrauding them of their money or personal information. As such, different scam types frequently overlap, blurring the lines between distinct categories and demonstrating the intricate web of fraudulent activities that exist today. The multifaceted nature of these scams highlights the difficulty in creating a comprehensive classification system that can effectively encompass all types of fraudulent schemes.
1.2 AI techniques
This study investigates AI-based techniques for processing unstructured text data to analyse fraud. Much of this text data, such as news articles, research papers, government reports, books, social media posts (such as tweets and Facebook comments), communications (such as emails, SMS messages, and chat logs), and web content (such as reviews on online marketplaces, travel and hospitality platforms, and comments on video sharing platforms), is inherently unstructured. Statista reported the global data created, captured, copied, and consumed is 64.2 zettabytes in 2020, and it is expected to exceed 180 zettabytes by 2025 Statista . With each new digital platform or communication channel, this data is increasing. Most of this created data is unstructured text data that provides opportunities for understanding human behaviour, habits, opinions and experiences. It contains information about users’ experiences, events, themes, opinions and sentiments that can be important for deriving meaningful insight from their experience related to fraudulent activities. Manual traditional data analysis techniques like keyword searches and coding of themes, are often limited and unachievable to extract meaningful insights, making advanced computer-driven automated techniques necessary.
However, this data often demonstrates significant challenges due to the diversity of natural (human) language. This includes dealing with noise, which includes irrelevant data, a wide array of linguistic variations of human language due to regional or cultural nuances, use of slang or jargon, abbreviations, spelling errors, typos and grammatical mistakes, which often pose challenges for analysing text data efficiently. NLP techniques were designed to effectively understand the structure (syntax) and comprehend (semantic) spoken and written human language the way humans do. Advancements in AI, including machine and deep learning, along with improvements in technology (such as increased computing power) and software (such as the availability of programming tools and libraries), have significantly improved the ability to process and understand large volumes of unstructured text. These tools have been widely used in many fields from sales and marketing to spam detection. The process of collecting, pre-processing, and training AI-based models using text data often involves the pipeline shown in Figure1:
Figure 1: Common pipeline for NLP-based models
Figure 2: Confusion Matrix
  • Problem statement: Using domain knowledge, a suitable research question is formulated for AI to address. This could be a classification problem (e.g. to classify text into a number of categories), or explanatory analysis involving identifying patterns within a text.
  • Data preparation: AI-based models require the collection of appropriate data to train models to analyse the research question. Often, the acquired data comes as unstructured data and requires cleaning and pre-processing. The data collection involves the collection of appropriate data and building a corpus. A corpus is a collection of large and structured sets of documents (e.g. emails, news articles, social media posts or transcripts). The data cleaning and pre-processing involves removing unwanted or redundant data to reduce the noise in the data. This may include removing duplicates or incomplete entries, removing symbols, punctuation, numbers, stop-words, converting acronyms to full words, and handling non-English words and slang or a particular jargon. Further pre-processing may involve text normalisation techniques like stemming or lemmatisation to reduce words to their root or base form to improve the accuracy of text analysis.
  • Feature engineering and selection: Feature engineering involves preparing data for machine learning models, and it involves the extraction and selection of predictive features in supervised learning) or find patterns and structures in unlabeled data in unsupervised learning. This task requires using domain knowledge to develop and select appropriate features. Common text features often used are n-grams (sequences of n consecutive words); Term Frequency-Inverse Document Frequency (TF-IDF) (a statistical method that weighs the importance of a word/term to a document within a corpus) matrix; sentiments and emotions present; lexical features (e.g. presence of certain words, Keyword-in-Context and lexical diversity); syntactic features (e.g. part-of-speech tags); semantic features (e.g. entities mentioned, word-embedding); readability scores; structural features (e.g. length of the text, number of paragraphs); and domain-specific features (e.g. presence of certain specialised terms).
  • AI technique/algorithm selection: This step involves selecting an appropriate AI algorithm for building the AI-based model. The text tasks often involve two main categories of AI-based models: supervised and unsupervised machine learning. The choice of the algorithm will depend on the learning, and type of AI required. Supervised machine learning algorithms are often used for text classification problems. The learning algorithm is provided with input features (training data) and labels (discrete outputs). The goal of the supervised machine learning algorithm is to map input features to discreet outputs. Traditional supervised machine learning algorithms include Logistic Regression, Naive Bayes, decision trees, Random Forest (RF) (multiple decision trees), Support Vector Machines (SVM) and K-Nearest Neighbors (KNN). New techniques of supervised machine learning include neural networks and deep learning-based models. Unsupervised machine learning models, often used in exploratory data analysis, involve working with unlabelled data to discover hidden patterns and themes. Unsupervised machine learning algorithms including clustering techniques using algorithms like K-means, hierarchical clustering and DBSCAN (Density-Based Spatial Clustering); and topic modelling using algorithms like Latent Dirichlet Allocation and Latent Semantic Analysis (LSA).
  • Model training: Often machine learning algorithms will have parameters that need to be tuned before learning begins, these are known as hyperparameters.The tuning process involves re-training the model multiple times using different values for these hyperparameters and selecting the best combination of values based on model performance on a metric of interest. In the case of supervised machine learning, the hyperparameters might be tuned using model performance on different “folds” of the data in an approach known as cross-validation, with some randomly-selected proportion of the data kept separate from the training data, known as a test dataset, for final model evaluation, giving the best available indication of how the model is likely to perform on new, unseen data. In the case of unsupervised modelling, heuristics will be used to identify the optimal number of clusters or topic modelling.
  • Model evaluation: The performance of the model needs to be evaluated. In the case of supervised modelling, this will involve measuring the performance of the model on the test data. The classic supervised machine learning algorithms can be evaluated using performance metrics such as confusion matrix (Figure 2), accuracy, precision, recall, F1-score, sensitivity, specificity, Receiver Operating Characteristic (ROC) curve, and the Area Under the Curve (AUC) curve:
    $Accuracy=\frac{TP+TN}{TP+TN+FP+FN}$
    $Precision=\frac{TP}{TP+FP}$
    $Recall=\frac{TP}{TP+FN}$
    $F1-score=\frac{2*Precision*Recall}{Precision+Recall}$
    Where:
    True Positives (TP): The model correctly predicts positive class (e.g., those that were classified as scam.)
    True Negatives (TN): The model correctly predicts negative class (e.g., those classified as not-scam).
    False Positives (FP): The model incorrectly predicts the positive class but it is not (e.g. not-scam is predicted as a scam).
    False Negatives (FN): The model incorrectly predicts the negative class but it is not (e.g. scam predicted as not scam).
    Sensitivity and specificity are two other performance measures. Sensitivity is the same as recall, or true positive rate that captures the model’s ability to correctly identify positive class (i.e. scam cases):
    $Sensitivity(TPR)=Recall=\frac{TP}{TP+FN}$
    Specificity, also known as false positive rate (FPR) measures the proportion of true negatives, and it captures the model’s ability to correctly identify negative class (i.e. not-scam cases):
    $Specificity=\frac{TN}{FP+TN}$
    The ROC curve is a graph illustrating the performance of one or more binary classifiers. It plots the sensitivity against the 1 - 1-specificity for various thresholds. The AUC is calculated as the area under the ROC curve.
  • Deploy model: Once the models are working well they can be deployed. When considering deployment of the model, one must address questions regarding why others should trust the model, how the model arrived at its conclusions and usability, and carefully assess the ethical implications of AI to ensure its suitability for deployment and that it is not biased.
In cases of unsupervised machine learning models, due to a lack of ground truth labels, the performance of the model evaluation may involve subjective interpretation to interpret the outputs (e.g. clusters or topics) generated by the model.
1.3 Advanced NLP techniques
This section provides a brief introduction to advanced NLP techniques, aiming to familiarise the reader with these concepts as they are later referred to during the SLR findings.
Word embeddings is a technique in NLP that encodes words as vectors of real numbers to capture their similarities. Words that are closer together in the vector space are expected to have similar meanings or relationships. Two of the widely used word embedding techniques are Word2Vec and GloVe. Word2Vec uses a simple neural network trained on large text datasets iteratively to predict either context words or target words. Word2Vec uses two approaches DBLP:journals/corr/Rong14 : Continuous Bag of Words (CBOW) predicts the target word based on its surrounding context words, whereas Skip-gram predicts surrounding context words based on a given target word. In the sentence ’The quick brown fox jumps over the lazy dog’, if ’fox’ is used as the target word, the CBOW model uses ’The’, ’quick’, ’brown’, ’jumps’, ’over’, ’the’, ’lazy’, and ’dog’ as context and predicts the word ’fox’. In Skip-gram, ’fox’ is used to predict the surrounding words like ’The’, ’quick’, ’brown’, ’jumps’, ’over’, ’the’, ’lazy’, and ’dog’. Global Vectors for Word Representation (GloVe glove learns the vector representation of words using global word-word co-occurrence statistics obtained from the training data to show the semantic relationships between words.
Word embeddings are often used by Large Language Models to comprehend and respond to language.
Large Language Models(LLMs) zhao2023surveylargelanguagemodels are advanced NLP tools trained on billions of words from a wide variety of sources such as newspapers, books and websites, which are designed to perform complex tasks such as translations, summarisation and the performance of human-like conversational abilities. Most LLMs are developed using a transformer-based architecture (transformers) DBLP:journals/corr/VaswaniSPUJGKP17 and they have billions of parameters used for training. Transformers are a type of deep-learning neural network model, and they are more efficient compared with predecessor state-of-the-art models which were based on Recurrent Neural Network (RNN). Transformers use a complicated architecture that has encoder and decoder layers to understand sequences of words and provide an output DBLP:journals/corr/VaswaniSPUJGKP17 . While the encoder layer processes input text data, extracting hierarchical representations through mechanisms like self-attention, the decoder layers generate output sequences based on the input received from the encoder. Transformer-based LLMs include GPT models including GPT3 (Generative Pre-trained Transformer 3), GPT-4 and GPT-4o by OpenAI. Both GPT-4 and GPT-4o are multimodal models that accept text and image and produce text chatgpt2024 . Other transformers include Bidirectional Encoder Representations from Transformers (BERT), and its smaller and lighter version of DistilBERT which is designed for applications where computational resources might be limited. BERT and DistilBERT are also designed to understand context in language processing and are suitable for NLP tasks like text classification, answering questions and named entity recognition. The difference between models like BERT and GPT is the way their architecture is designed and the objectives of learning. BERT uses only the encoder of the transformer architecture and processes text bidirectionally by considering both preceding and following words. On the other hand, GPT models use only a decoder of the transformer architecture and process text in an unidirectional manner by processing text based only on preceding words.
LLMs can assist in analysing large amounts of text data and identify patterns automatically which can be useful when dealing with fraud and other crime-related data, they can be misused by criminals to generate content for fraudulent activities. Generative AI refers to AI techniques that can be used to generate new text, audio, images and video that closely resembles human-generated content. LLMs have been successfully applied in various areas of human communications including chatbots in customer support systems by generating human-like text, content generation and performing language translation. However, this same technology can be exploited by criminals to create content for deception including fake websites, targeted phishing emails and scam advertisements to deceive potential victims.
1.4 Research approach
Although there are some literature reviews related to the application of AI for fraud and crime, to our best knowledge there are currently no SLRs that attempt to understand the state-of-the-art towards detecting online fraud in general. Overall, the literature reviews that we found discuss the state-of-the art towards detecting specific online fraud or scams, e.g., credit card fraud alone. In this work, we aim to understand whether there are universal AI methodologies that attempt to detect online fraud, focusing on textual data.
Search String
("Online Fraud*" OR "scam*") AND (("machine learning")
OR "NLP" OR ("natural language processing") OR "classifier"
OR ("Large Language Models") OR "LLM"
OR ("Generative Artificial Intelligence") OR "GenAI" OR "GAI")
Table 1: Query for the literature selection in various academic libraries.
2 Systematic Review
Systematic reviews differ from traditional literature reviews as they aim to identify all relevant studies that address a set of research questions using a methodology that is structured and can be replicated nightingale2009guide .
2.1 Methods
The following methodology was employed to conduct the SLR, and address the selection process for identifying relevant publications and to avoid biases.
2.1.1 Protocol
For this SLR we follow the Preferred Reporting Items for Systematic Reviews and Meta-Analysis extensions for Scoping Reviews (PRISMA-ScR), as proposed by Moher et al. moher2010preferred .
2.1.2 Eligibility
This review focuses on studies that use AI-based models, specifically NLP models, including Machine Learning (ML) and Deep Learning (DL) techniques using text data. For example, studies that employ AI to detect fraudulent bank accounts, fraudulent credit card transactions, or fraudulent networks of users online, are not to be included.
To be considered for inclusion in the SLR, we used the following eligibility criteria:
  • Peer-Reviewed Studies: We focused on peer-reviewed studies published in English between 2019 and March 2024. Our search was restricted to academic records found in journals and conference proceedings. We excluded theses, legal documents, patents, and citations from these studies.
  • Grey Literature:To capture the latest AI-based models, we also included grey literature, specifically pre-prints from ArXiv published in 2023 that have not yet been incorporated into conference proceedings or book chapters.
  • Search Strategy: Table 1 shows the search strings used to detect related literature in ACM Library, ProQuest, Web of Science, IEEE Xplore, ArXiv, and Google Scholar. The authors finalized this string after trying various searches in these libraries.
  • Scope and Focus: Studies must address fraud performed online and use AI-based methodologies for analysing online fraudulent activities. The focus is on studies using textual data, whether from scammers, victims, or victim reports, to understand, detect, or analyse online fraud activities. Our goal is to understand what is the state-of-the-art on models that are designed to prevent and detecting scams (tricking the victim) before the victim gets defrauded which will lead them to monetary losses. Studies that analyse money transactions, credit card transactions, and cryptocurrency transactions are excluded, from this review, as they do not use text data.
  • AI-Methodology: Papers must include a methodology section or a similar section where the authors discuss their AI implementation and fine-tuning along with the accuracy of their model. Finally, we consider studies published after 2019.
  • Publication timeframe: Papers published between 2019 and March 2024 are included in this review. This period was selected to manage the overwhelmingly large volume of papers on online fraud and to align with our available resources. Also, we believe that studies conducted before 2019 are less likely to reflect recent advancements in AI methods. Given the rapid evolution of AI-based models, focusing on the 2019-March 2024 period ensures the inclusion of the most up-to-date and relevant research.
Label Description Example
Title The title of the manuscript Detecting Phishing URLs Using Deep Learning
Author Author’s full name plus the abbreviation et al. if applicable Smith, John or Smith et al.
Year The year the work got published (YYYY) 2020
Fraud Type The type of scam the authors try to detect, analyse, or discuss in their manuscript Phishing URLs
Data Type The type of data the authors use for their analysis URLs
Data Quantity The amount of data used for the analysis 5000 phishing URLs and 5000 legitimate URLs
Models Used All models the authors experimented with Random Forest, Linear Regression, LDA, W2V
Best Model The model with the best accuracy Random Forest
Model Stats All performance metrics A=0.95, P=0.80, R=0.9, AUC=0.89
Table 2: Data items and characteristics extracted from the literature.
2.1.3 Data extraction
Next, the authors agreed on the data that should be extracted from the included studies. The task of extracting data was carried out by only one of the three reviewers. This was deemed sufficient since the reviewer’s role involved extracting the required details from the papers, and it was unnecessary for a second reviewer to check accuracy. The only aspect of the data extraction that the reviewer had to conceptualize was the specific Fraud Type analyzed by the study under question. For example, if a study is analyzing Phishing URLs, then it is labelled as Phishing URLs.
Not all papers explicitly specified the type of fraud they were analysing. Due to the diversity of scams, there is no agreed way of labelling fraud types. As as result, we used an umbrella term to categorise them. For example, online scam campaigns, made by bot users to advertise fraudulent phishing URLs include fake users and phishing URLs fraud analysis, hence we agreed to label papers of broad online scam campaigns as social media scams.
To ease the representation of our findings, we do not classify a paper with more than one fraud type. More specifically, the reviewer had to resort to the main goal of the paper towards annotating the kind of fraud it analyses. For example, if a paper used phishing emails to extract phishing URLs towards detecting phishing URLs, then that paper would be labelled as Phishing URLs, as that is the main goal of the study. The final version of the data extracted from each record is depicted in Table 2, along with relevant examples.
3 Results
3.1 Study selection and characteristics
The PRISMA-ScR flow diagram in Figure 3 summarises the study selection process for the two academic searches. $2,617$ stuides were identified for eligibility screening. The ACM Digital Library returned $389$ documents, IEEE Xplore returned $712$ documents, Web of Science returned $253$ documents, ProQuest returned $783$ documents, Google Scholar returned $399$ documents, and ArXiv returned $47$ documents. An additional $15$ papers were recommended by the experts in the area. After removing duplicates, $2,457$ papers remained for further review.
Figure 3: PRISMA Chart
At this stage, $10\%$ of these papers were selected ($N=242$) for Inter-Rater Reliability, to calculate the multi-annotator agreement between the three annotators of this review. The Fleiss Kappa score between all three annotators was $0.83$, indicating almost perfect agreement. The Cohen Kappa Agreement was also calculated between each pair of annotators. The agreement between the annotator A and the annotator B was $0.65$ (substantial agreement), between A and annotator C was $0.66$, and between B and C was $0.52$ (moderate agreement). [^1] Annotators A, B, and C will be replaced with the initials of authors after the reviewing process. The three annotators compared their annotation process and reviewed the eligibility criteria and goals of this review. Following this discussion, the lead annotator moved to perform the rest of the annotation of the papers to be included in this review.
Reviewing the abstract of those papers resulted in $2,107$ papers being excluded from the study as they did not fit the eligibility criteria discussed in Section 2.1.2. This resulted in $350$ full-papers passing to eligibility screening, out of which $127$ did not fit the eligibility criteria and was excluded. Overall, this process resulted in $223$ full-text papers being included for qualitative analysis.
Figure 4: Percentage of scam types analyzed in the studies included for qualitative analysis
3.2 Types of Online Fraud
Figure 4 shows the types of fraud analyzed in the full-text papers that were included in the qualitative analysis. The reviewer of these studies manually coded each paper with the scam type that the study is focusing on, based on the title, abstract, and methodology of the studies. The majority of studies that were included for qualitative analysis focus on phishing detection, with about a third ($29\%$) of the studies focusing on detecting phishing URLs online ($N=64$). More specifically, these works tackle the problem of automatically detecting whether a given URL is likely to be fraudulent. A large number of papers were related to detecting phishing emails ($N=29$), followed by studies on SMS phishing detection ($N=20$). Other studies on phishing include phone call transcripts towards understanding and detecting voice call phishing ($N=12$) and a small number of studies attempt to understand phishing methods via victim reports ($N=4$).
Moving on to other types of fraud, we found many studies that attempted to detect fake reviews on various platforms like Google Play Store, Apple App Store, Yelp, and TripAdvisor ($N=23$). Another widely studied scam focus was recruitment fraud ($N=20$). We also found a number of studies that employed AI techniques to detect fake accounts on Facebook, Instagram, and Twitter ($N=18$).
A few studies used Generative AI (GenAI) to gain a better understanding of some of the latest methods of phishing and other fraud types. GenAI represents cases where advanced LLMs are misused towards social engineering attacks. GenAI models, have revolutionised various industries as they are now able to promptly generate human-like text in response to input commands. $3$ studies employed GenAI models to automatically interact with scammers to waste the scammers’ resources or time, and to collect information regarding the methods with which the scammers try to defraud users to disrupt their operation. This is defined as scam baiting, the process of using generative AI models to deceive and engage with scammers. Interestingly, this is a countermeasure against online fraud (GenAI Scam Bait in the figure).
Our search also returned many studies that discuss the misuses of GenAI towards Social Engineering Attacks ($N=13$), where scammers use these advanced models to create legitimate-looking emails or SMS to earn the trust of potential victims. While GenAI models offer numerous benefits, these studies show that they pose significant risks when leveraged for malicious purposes, particularly in the realm of social engineering. GenAI can generate coherent, contextually relevant, and grammatically correct emails that mimic the style and tone of professional communication. This increases the likelihood of victims perceiving fraudulent emails as legitimate and trusting the message schmitt2023digital .
Regarding other scams, we found a variety of studies that try to detect Social Media Scams ($N=6$). These scams included a variety of fraudulent activities including fake user accounts and online groups, advertisement of fraudulent apps or fake phishing website aimed at stealing personal information or money.
Similarly, $3$ studies focus on Romance Fraud via analysing profiles on social media and discussions with victims. We also identified $3$ studies that try to understand fraudulent cryptocurrency investment scams and $2$ studies that attempt to detect the likelihood of cryptocurrency manipulation. Finally, we identified $2$ studies that analysed fraudulent crowdfunding online and $1$ that studies fraudulent e-commerce websites.
4 Summary of Findings
We now provide a summary of our findings, categorized per fraud activity analyzed within the papers that are included in our SLR for qualitative analysis.
4.1 Data Sources
First, we report the most popular data sources used and the datasets analyzed.
Phishing URLs. We start with understanding the preferable data sources and methodologies employed for the analysis and detection of Phishing URLs, as it is the most popular scam type category we detected in our SLR. In total, we analyze the data sources and detection methodologies of the identified $63$ papers that attempt to tackle this issue. Table LABEL:tbl:data_phishing_URL summarizes the data sources and methodologies used to detect Phishing URLs, as found in the literature.
First, we find that researchers used various websites that offer information on URLs for the analysis of malicious and legitimate domains. This information may be webpage rankings (how trusted the webpage is), phishing reports, and historical data. By far, the most popular data source used was PhishTank, [^2] https://phishtank.org/ a website that allows users to report webpages that might be malevolent or suspicious, with $25$ studies using it as already labeled malicious websites li2022mui ; vo2023shark ; al2020convolutional ; aslam2023phish ; jishnu2023enhanced ; jaber2022improving ; rafsanjani2023qsecr ; alswailem2019detecting ; mandadi2022phishing ; jha2023intelligent ; marimuthu2022intelligent ; adebowale2023intelligent ; shaiba2022hunger ; rao2020two ; salloum2021phishing ; orunsolu2022predictive ; rao2019detection ; barraclough2021intelligent ; almseidin2019phishing ; chiew2019new ; li2019stacking ; sahingoz2019machine ; somesha2020efficient ; tharani2020understanding ; do2020malicious ; pradeepa2022lightweight .
Another website that offers a list of phishing URLs is OpenPhish [^3] https://openphish.com/ and it was used by $3$ studies vo2023shark ; almseidin2019phishing ; chiew2019new . Two studies used URLhaus [^4] https://urlhaus.abuse.ch/, a project for sharing malicious URLs, for the collection and analysis of phishing URLs rafsanjani2023qsecr ; do2020malicious . We also find one work that used SpamHaus [^5] https://www.spamhaus.org/ for the collection of IP and domain reputation fernandez2022early , and one that used URLscan [^6] https://urlscan.io/ chen2021ai . Interestingly, chen2021ai also collected user-reported domains from ScammerInfo [^7] https://scammer.info/, a forum where users post and discuss various scams. Lastly, the webpage WhoIs [^8] https://who.is/, a webpage that offers historical data on webpages, was used for feature collection from shalke2022social ; adebowale2023intelligent .
The most used data source for collection of legitimate webpages was Amazon Alexa, a webpage that offers webpage rankings, with $10$ studies using it to collect legitimate annotated webpages li2022mui ; navyah2022ensemble ; saha2023phishing ; liang2019using ; shaiba2022hunger ; orunsolu2022predictive ; rao2019detection ; somesha2020efficient ; tharani2020understanding ; do2020malicious ; liang2019using . Google’s Search Engine was used for one study chen2022development , while another used Majestic Million site [^9] https://majestic.com/reports/majestic-million for legitimate webpage collection jishnu2023enhanced , a site similar to Alexa.
We also find that many records use previously publicly available datasets for their analysis. More specifically, $11$ studies gu2022ensemble ; jha2023machine ; jha2023machine ; mehndiratta2023malicious ; jain2023support ; saha2020phishing ; kumar2023hybrid ; ashwitha2023perception ; dr2023malicious ; pradeepa2022lightweight ; zamir2020phishing used already publicly available datasets published on Kaggle. Similarly, other $5$ studies navyah2022ensemble ; jaber2022improving ; kalabarige2023boosting ; mohammed2022accuracy ; priya2020gravitational used the UCI Phishing Dataset. [^10] https://archive.ics.uci.edu/dataset/327/phishing+websites Other studies used publicly available datasets from other sources ariyadasa2022phishrepo ; villanueva2022application ; vecile2022malicious ; kumar2020phishing ; zin2023machine ; pradeepa2022lightweight ; pathak2023classification .
The most recent studies (published in 2023) that attempt to detect Phishing URLs automatically, attempt to do so by collecting data from alternative sources like social networks, and user-reported phishing URLs bitaab2023beyond ; saha2023phishing ; saha2023phishing ; bitaab2023beyond ; nakano2023canary ; janet2022real , while others used datasets from Telecom and Security institutions mehndiratta2023malicious ; yu2024efficient ; liang2019using Alas, we failed to detect the data source used for $9$ studies, as they do not clearly report how or from where they acquired the dataset they use for their study alkawaz2022identification ; el2021malweb ; kundra2023identification ; chen2020intelligent ; ou2021no ; raja2021lexical ; yadollahi2019adaptive ; nagy2023phishing ; geyik2021detection .
Phishing Emails. We now discuss the data sources used in the $29$ works that tackle the Phishing Email detection problem. The data extracted from the literature and presented in this section are depicted in Table LABEL:tbl:data_phishing_emails.
Two works genc2021understanding ; jiang2024detecting used emails received on the author’s personal or professional email spam folder. Other works that include datasets from alternative sources is the one of Mehdi et al. mehdi2023adversarial , which used various techniques to develop their dataset, like GPT2 generated synthetic phishing emails made available by previous research radford2019language , along with TextAttack [^11] https://github.com/QData/TextAttack, a Python framework for adversarial attacks, data augmentation, and model training in NLP, Textfooler [^12] https://github.com/jind11/TextFooler, a Model for Natural Language Attack on Text Classification, and PWWS ren2019generating . Another alternative data source for phishing email detection was used by Janez et al. janez2021trustworthiness who used data from SPAM Archive [^13] http://untroubled.org/spam/, a website that publishes spammy email repositories at the end of every month and is constantly updated.  gallo2019identifying used user reported e-mails. Lastly, the data source used by three works was not clearly stated within the manuscript mughaid2022intelligent ; venugopal2022detection ; rahmad2020performance .
Phishing SMS. Regarding Phishing SMS (smishing), we included and analyzed $20$ papers in this SLR. For the detailed data, refer to Table LABEL:tbl:data_smishing.
Similarly to previous analyses, we find that the overwhelming majority of works opted for using already publicly available datasets for their analysis and training a model to detect fraudulent SMS automatically. More specifically, a variety of subsets from a publicly available dataset on Kaggle [^14] https://www.kaggle.com/datasets/uciml/sms-spam-collection-dataset was used by $14$ studies vinothkumar2022detection ; agrawal2023effective ; jain2019feature ; mishra2023dsmishsms ; abid2022spam ; kohilan2023machine ; siagian2023improving ; gandhi2023sms ; al2023multi ; zhang2023bert ; dharani2023spam ; addanki2023safeguarding ; ulfath2021hybrid ; jain2022sms . Another study zhang2020lies used the Kaggle dataset, but incorporated Fake Base Station data and made it available to researchers. [^15] https://github.com/Cypher-Z/FBS_SMS_Dataset Similarly, this work vinothkumar2022detection used a subset of the Kaggle dataset in combination with emails and YouTube comments for spam content detection, while Lai et al. lai2022semorph used data provided by users. [^16] https://www.datafountain.cn/competitions/508 Tang et al. tang2022clues collected tweets where users were reporting smishing for their analysis. Two other works used data from the Korean Internet and Security Agency seo2024device and 360 Mobile Safe liu2021detecting . Lastly, Timko et al. timko2023commercial proposed a platform where users can freely post Phishing SMS for researchers to use. [^17] https://smishtank.com/
Phishing Phone Calls. We identified $12$ studies that used phone call transcripts to understand Voice call-enabled phishing (vishing).
Derakhshan et al. derakhshan2021detecting used the CallHome dataset which includes 120 unscripted 30-minute telephone conversations between native speakers of English. [^18] https://catalog.ldc.upenn.edu/LDC97S42 Another work djire2023evaluating used AI-generated deepfake voice recordings (Tacotron 2191919https://pytorch.org/hub/nvidia_deeplearningexamples_tacotron2/, Deepvoice 3202020https://r9y9.github.io/deepvoice3_pytorch/, and FastSpeech 2 ren2020fastspeech ). For authentic voice recordings they used the synplaflex dataset sini2018synpaflex which is a corpus of 87 hours of audiobooks in French.
Various works used telecommunication operator datasets, like liang2023telecom using fraudulent caller IDs and phone transcripts huang2022learning from telecommunication operators in China, hu2023urf4cct used data from the Public Security Bureau in Zhejiang Province, China, and kim2021voice used data from the Korean Financial Supervisory Service. Kale et al. kale2021classification developed their dataset via questionnaires and victim testimonies. Other works collected data from various social networks, like YouTube transcripts rahman2022phone , Facebook, online blogs and forums, public datasets, as well as some that were developed based on studies of scammers’ activities and behaviors hong2023scam . Others opted for using previously analyzed and publicly available data gowri2021detection ; malhotra2023detection , while the data used from Zhong et al. zhong2020encoding was unclear.
User Reports. Four studies used user reports to understand phishing activities.
First, liu2021fine constructed a fraud complaint dataset from the Internet finance service in China. Similarly, zhou2022keyword used court documents from Chinese online judgement records, while palad2019document used incident record forms from victims and interviews in the Philippines. Lastly, lwin2023supporting launch and introduce a website operated by the National Crime Prevention Council (NCPC) in Singapore, where users can report and get informed on the latest phishing activities.  [^21] https://www.scamalert.sg/
Fake Reviews. We move to the Fake/Fraudulent Review scam type. We include $23$ studies in our analysis and the data extracted from each record in our analysis is depicted in Table LABEL:tbl:data_fake_reviews
Other studies used application reviews from Google Play Store or Apple’s App Store yugeshwaran2022rank ; tushev2022domain ; obie2022violation . Other’s used previously available Amazon product reviews, [^24] https://snap.stanford.edu/data/web-Amazon.html or collected Amazon reviews rangar2022machine ; iqbal2023efficient ; furia2020tool ; gupta2020detecting ; thilagavathy2023fake ; chandana2021analyzing ; akshara2023small ; deekshan2022detection .
Others collected Amazon Hotel and Holiday packages reviews rangari2022empirical ; silpa2023detection , or TripAdvisor review data tufail2022effect Lastly, bevendorff2024product used YouTube transcripts to interpret false review exaggeration. The source of the data used by Ganesh et al. ganesh2023implementation was unclear.
Fraudulent Recruitment. We detected $19$ studies that focus on the detection of fraudulent job postings. The related data is listed in Table LABEL:tbl:data_recruitment.
The overwhelming majority ($N=16$), of papers used the same publicly available dataset from Kaggle, [^25] https://www.kaggle.com/datasets/amruthjithrajvr/recruitment-scam, which holds about 18K job postings out of which 800 are fraudulent. Notably, this dataset includes data from 2012 to 2014 nessa2022recruitment ; prathaban2022verification ; pandey2022effective ; ranparia2020fake ; habiba2021comparative ; reddy2023web ; santhiya2023fake ; lal2019orfdetector ; nasser2021online ; amaar2022detection ; sofy2023intelligent ; vo2021dealing ; nanath2023investigation ; ullah2023smart ; bhatia2022detection ; li2021exploratory .
The other three studies developed custom crawlers to collect data from various job posting sites in the UK (SEEK, Glassdoor, Indeed, and Gumtree) mahbub2022online , in Bangladesh (job.com.bd, bdjobstoday, deshijob) tabassum2021detecting , and in China (China–Boss, Zhipin, Liepin, and 51job) zhang2023orfpprediction .
Fake Accounts Online. Various studies attempt to tackle the automated detection of fake profiles online. In Table LABEL:tbl:data_fake_users we depict the data extracted from the identified records.
We find that many works collect user profile data from various online social networks for this analysis, like Twitter raj2020multi ; gangan2023detection ; yue2019madafe ; singh2023safe ; ali2019detect ; shukla2022tweezbot ; rovito2022evolutionary , Instagram das2022effecient ; fathima2023ann ; anklesaria2021survey , Facebook albayati2019identifying ; venkatesan2019graph ; shreya2022identification , YouTube na2023evolving , and Sina Weibo zhang2021social . A different approach was used by a study haq2023spammy that collected real names from various webpages, schools, and other sources to automatically detect fake names online.
Other works used previously published and openly accessible datasets that included user data from various social networks nikhitha2023fake ; bebensee2021leveraging
GenAI-facilitated Social Engineering Attacks. Under this category, we find many works that investigate how generative AI models can be misused to defraud people.
Various studies janjeva2023rapid ; schmitt2023digital ; ferrara2024genai develop and discuss an initial taxonomy where they present how AI-generated content can be misused by scammers. At the same time, Carlini et al. carlini2021extracting test various membership inference attacks on OpenAI’s GPT2 model and confirm that the model is vulnerable to this kind of attack which poses risks of privacy. Similarly, Kumar et al. kumar2023augmenting discuss the significant implications for cybersecurity, privacy, and ethical considerations that should be considered when developing and using these models.
Moving on to misuse cases of these models, ayoobi2023looming discuss how LLMs and GenAI can be used for fake professional profile bios to trick users that the account is legitimate. Similarly, diresta2024spammers show that scammers can use these models to create AI-generated images to be posted on various social networks. Their case study shows that these images tend to receive high volumes of engagement on Facebook as many users do not seem to recognize that the images are synthetic. Other works show how these models can be jailbroken to produce code to immitate legitimate webpages (phishing URLs) grbic2023social , malware code, phishing emails, phishing SMS, SQL injection attacks, and many more potentially dangerous material shibli2024abusegpt ; alotaibi2024cyberattacks ; alawida2024unveiling .
At the same time other works show that humans may be able to accurately detect phishing AI-generated content sharma2023well , while Roy et al. roy2023chatbots discuss and experiment with countermeasure to prevent malicious prompts (jailbreaking) in GPT and they provide valuable insights into how the model can become more robust against this vulnerability.
Social Media Scams. Six studies examine various scams and spammers facilitated through various Social Networks. xu2022efficiently used data from WeChat and Konect repository to detect users that use WeChat to defraud people. la2023sa and la2021uncovering used Telegram data to characterize and detect Fake Telegram channels, while  shah2020illicit collected data from Telegram and compared it to Twitter data to understand and detect fake users. Similarly, al2023dspamonto collect and analyzed Twitter and Institute of Informatics and Telematics data to detect scammers on Twitter. tripathi2022analyzing collect Youtube data to detect scammers that attempt to lure in victims via YouTube comments on videos.
GenAI Scam Baiting. We identified three works that used LLMs and Generative AI to automatically engage with scammers online towards wasting their resources and collect data on various fraud activities. For this section we do not identify data sources used as researchers use data received on their baiting accounts.
Romance Fraud. Moving on to Romance Fraud, he2021datingsec attempt to automatically detect malicious accounts on Momo, a dating website. Similarly, suarez2019automatically collect data from DatingMore and scamdigger.com to develop automated methods to understand fraudulent profiles within dating social networks. Lastly, lokanan2023tinder analyzed the sentiment of tweets with the hashtag #tinderswindler to provide an understanding on users sharing their experiences regarding romance fraud.
Fraudulent Investment. Studies that attempted to understand fraudulent investment scams did so employing various datasets and methodologies.
First, siu2022invest analyze investment scam advertisements found in Bitcointalk. [^26] https://bitcointalk.org/ Li et al. li2023towards collect YouTube comments to automatically detect bots that advertise fraudulent investment content. Lastly, kuo2024constructing develop a scam detection model based on emotional fluctuations of user discussions collected from one of the most popular instant messaging app in Taiwan.
Crypto Manipulation. Market, and more specifically, cryptocurrency coin manipulation is when users collectively attempt to alter investor interactions towards manipulating the prize of a coin.
nizzoli2020charting discuss this process in detail via data acquired from Twitter, Telegram, and Discord channels. Similarly, mirtaheri2021identifying identify and analyze cryptocurrency manipulations from user activity collected from Telegram and Twitter.
Fraudulent Crowdfunding. Both works that analyse fraudulent crowdfunding lee2022backers ; shafqat2019topic collect the descriptions and metadata from hundreds of Kickstarter campaigns. [^27] https://www.kickstarter.com/
Fraudulent E-commerce. salihovic2019role analyze the Terms and Conditions of various websites that sell various products towards understanding and detecting obscured financial obligations in online agreement.
4.2 Methodologies employed
We now discuss the most popular AI and NLP methodologies employed per online frayd type.
Phishing URLs. We find that the records included in our SLR attempt to automatically detect phishing URLs using a variety of NLP and AI methodologies. These models included classic supervised machine learning algorithms such as Naive Bayes (NB), Random Forest (RF), Decision Tree (DT), Support Vector Machines (SVM), XGBoost (Extreme Gradient Boosting), KNN (k-Nearest Neighbors), as well as Artificial Neural Networks (ANN) and more advanced in deep learning such as Long Short-Term Memory (LSTM) and Convolutional Neural Network (CNN). LSTMs are a type of Recurrent Neural Network (RNN) designed to capture long-dependencies in sequential data, making them suitable for handling and predicting sequences of text. On the other hand, CNNs aim to identify key features in the text by capturing local patterns. These models were developed for binary classification tasks, whether a URL is fraudulent or not.
NLP techniques related to text mining have been used to extract features from URLs, which are then used as features to train the AI-based classification models. More specifically, counts of characters, special characters, n-grams, and URL deconstruction: detection of secure scheme or not, domain and top-level domain, along with sub-directories. Various works used hybrid, or combination of methodologies for their detection including some more advanced techniques, like li2022mui using Bi-LSTM (Bidirectional Long Short-Term Memory that could process sequences of text in both forward and backward directions) with VGG (Visual Geometry Group - a type of CNN), vo2023shark using W2V (Word2Vec) for feature extraction along with CNN, GRU (Gated Recurrent Unit, - like LSTM but a more simplified RNN), and Bi-LSTM, nakano2023canary using BERT with RF, bitaab2023beyond developed a hybrid system that uses RF, SVM, FNN, and XGBoost, alswailem2019detecting used Linear Regression (LR) and DT, and villanueva2022application used LSTM and GRU.
Other stand-alone AI methodologies, like LightGBM (Light Gradient Boosting Machine - an ensemble learning technique designed for handling large datasets with large features), RF, NB, and ANN also seem to work well on detecting phishing URLs, but by far the one with the best performance seems to be RF.
Phishing Emails. The methodologies used for phishing email recognition weigh more on the NLP analysis. The majority of studies used various NLP methodologies for feature extraction, including, but not limited to topic analysis (LDA, BERT, BERTLARGE), n-gram extraction (TF-IDF, BOW, Clustering, W2V), and sentiment analysis (VADER, WordNet).
Then, the studies that also opted for automated detection of phishing emails employed LLM analysis, RF, NB, SVM, CatBoost, LSTM, RNN, and many more.
Phishing SMS. Similar to phishing email detection, phishing SMS detection also relies on state-of-the-art NLP methodologies, including LLMs, LDA, BERT and W2V. The existing literature also used AI methodologies like LR, SVM, CNN, GNN, LSTM, NB, and KNN for automated detection.
Phishing Phone Calls. Studies on Vishing used various AI-methods for automated detection. The majority of studies used transcript text data for their analysis, with the exception of djire2023evaluating who analyzed deepfake voice analysis. For that study, the authors found that RNN had the best performance.
On text data, various NLP and AI methodologies were used, including but not limited to SVM, NB, LSTM, CNN, RF, BERT, W2V, LR, and KNN.
Phishing User Reports. We find that for the analysis of user reports of various phishing activities, the use of BERT, SMO, J48, NB, RF, XGBoost, D2V, Jaccard, NER, and TF-IDF were applied.
Fake Reviews. The most popular NLP methodologies that apply on Fake Review detection lean heavily on Sentiment Detection methodologies including VADER and WordNet. Similar to previous fraud analysis, the AI methods applied include various Neural Network Models like CNN.
Fraudulent Job Postings. The methodologies employed for the automated detection of Fraudulent Job Postings include stand-alone machine learning algoritunsm used for classification tasks including LR, SVM, KNN, RF, XBoost, and ANN, and deep learning models including Bi-LSTM.
Scam Baiting. One of the Scam Baiting studies,  cambiaso2023scamming used OpenAI’s ChatGPT to reply to scammer emails. Similarly, bajaj2023automatic experimented with OpenAI’s ChatGPT and DistillBERT to categorize various scam emails they received and provided a qualitative analysis of how well the two models perform. The other study, chen2023active set up a mailserver and used data from Scambaiter mailserver, Enron Email Dataset, and ScamLetters.Info while employing their own DistillBERT model to engage with scammers towards categorizing and analyzing the various emails.
Romance Fraud. Studies on Romance Fraud detection used various NLP methodologies, including sentiment detection using BOW, TF-IDF, and textBlob, finding that RF works best. For the detection of malicious accounts in dating applications, LSTM works best, while another study  suarez2019automatically found that EML also works well.
Fraudulent Investment. Three of the studies related to Fraudulent Investment used data from different sources (forum, messaging app, and YouTube). The reported studies indicate that DT was the best performed machine learning model for detecting emotional flactuations on victims, while XGBoost worked well on detecting malicious advertisement of fraudulent investment websites.
Crypto Manipulation. The two identified studies on Cryptocurrency Market Manipulation used pre-existing methods for detecting fake users, along with CorEx Topic Analysis nizzoli2020charting . The other study found that SVM with SGD and TF-IDF works best for detecting discussions that aim to manipulate the market mirtaheri2021identifying .
Fraudulent Crowdfunding. One of the papers on fraudulent crowdfunding detection used NLP methodologies including Named Entity Recognition and other NLP featurues detected in the descriptions of Kickstarter campaigns and build a LR model that performs well lee2022backers . The other study, shafqat2019topic developed an LSTM-LDA topic detection model that analyses the crowdfunding campaign, along with the comments of people towards estimating whether a campaign is a scam or not.
Fraudulent E-Commerce. Finally, the only study that we identified that used text data towards understanding Fraudulent E-Commerce activities, used OpenAI’s GPT-4 model to automatically detect obscure financial obligations in the Terms and Conditions of the focused websites salihovic2019role .
4.3 Key Findings
Given the findings of this SLR, we attempt to address our Research Questions.
Phishing URLs. While established datasets have played a crucial role in the development of phishing URL detection models, there is a clear need to incorporate more dynamic and current data sources. Leveraging user-reported phishing URLs from social networks, along with data from Telecom and Security organizations, offers a more effective approach to combating phishing attacks. These sources provide real-time, diverse, and relevant data that enhance the robustness and accuracy of detection models, keeping pace with the evolving nature of phishing threats. By combining the strengths of both traditional and modern data sources, researchers can develop more comprehensive and adaptive phishing detection systems, better protecting users from phishing URLs.
Regarding the methodologies used, we find that the existing literature used state-of-the-art methodologies for the analysis and detection of phishing URLs. Notably, RF seems to be the stand-alone model that works best, while other hybrid methodologies also report promising performance.
Phishing Emails. While publicly available datasets have laid the groundwork for phishing email detection research, the rapidly evolving nature of phishing attacks requires the use of more dynamic and up-to-date data sources. Leveraging user-reported emails, real-time spam collections, and advanced synthetic data generation techniques can significantly enhance the robustness and accuracy of phishing detection models. By combining traditional datasets with innovative data sources, researchers can develop more comprehensive and adaptive phishing detection systems, better equipped to detect phishing activities via email.
All of the works that opted for automated phishing email detection report very good performance on their detection models, with RF, BERT, LSTM, RNN, and SVM being the most popular.
Phishing SMS By this point we established that many works in various scam type analysis rely on established datasets. Alas, the rapidly evolving nature of smishing requires a more dynamic and diversified approach to data collection. By integrating publicly available datasets with real-time, user-reported data and specialized security sources, researchers can develop more effective and resilient smishing detection models. This approach will ensure that models remain relevant and capable of addressing new and sophisticated smishing threats as they arise.
Contrary to the phishing email detection, we find that in the case of phishing SMS detection, which tends to be much smaller text, SVM and various applications of Neural Networks perform best.
Phishing Phone Calls. Regarding the data sources used for automated vishing detection, most works used text data via transcribing voice recordings. Other works also used caller ID information from various telecommunication operators. Some works collected data from user reports and only one work attempted to detect deepfake signals in voice recordings. The best performing methodology used for automated vishing detection is SVM.
Phishing User Reports Reviewing the four works that studied phishing via user reports, we find that researchers used data from various sources, including court judgements, public data from forums, and user reports from Financial Institutions. The methodologies applied vary as there is use of Named Entity Recognition, various NLP methodologies, and AI methodologies.
Fake Reviews. Although many works used previously available datasets to establish and compare their detection models, we notice a clear trend where later studies tend to collect data from platforms like Amazon, Google Play, Apple App Store, and YouTube for their analysis. This is extremely encouraging as the data used for these detection models need to be constantly updated. Hybrid models including LR, SVM, CNN, and LSTM seem to perform best for the detection of fake reviews.
Fraudulent Job Postings. While the Kaggle dataset has been pivotal in advancing research on fraudulent job postings, the rapidly evolving nature of job scams necessitates the use of more current and diverse data sources. Custom data collection methods, which tap into active job posting sites, represent a crucial step forward in enhancing the effectiveness of detection models. By leveraging a mix of established and new data sources, researchers can develop more comprehensive systems to effectively combat fraudulent job postings. The models that perform best for this fraudulent activity vary. We find that Bi-LSTM, KNN, RF, DNN, and LightGBM perform well.
GenAI Social Engineering. Only two works here use data collected from real use cases, namely Ayoobi et al. ayoobi2023looming discuss fake profile AI-generated content on professional social networks, while Diresta et al. diresta2024spammers present deep fakes posted on Facebook. Other works discuss the potential sybersecurity, privacy, and ethical issues with these models and how they can be misused to automatically create fraudulent content.
5 Discussion
We now discuss our findings, detailing recognised limitations and shortcomings detected in the reporting of AI models related to the performance and data sources used. We also provide recommendations for researchers developing detection models for online fraud.
5.1 Data Sources
Overall, we analyzed the data sources and detection methodologies of 222 papers that aim to address a range of online fraud problems. Although our findings reveal a preference for well-established datasets, especially in the automated detection of various phishing and fake reviews detection, more recent studies (published after 2023) seem to shift towards more dynamic and recent sources.
We found that the overwhelming majority of the Phishing domain detection studies relied on well-known and extensively studied datasets from websites like PhishTank, OpenPhish, and SpamHaus. In contrast, others used datasets made available from previous studies or publicly available repositories like Kaggle, University of California Irvine (UCI) Machine Learning Repository, and GitHub. This is also the case for studies that focused on phishing email detection, fake review detection, and fraudulent recruitment detection. However, while these established datasets provide a valuable foundation for research, they come with limitations.
Online fraud is dynamic, with new scam techniques frequently emerging and older methods, like phishing URLs and websites, continually evolving. Studies show hat LLM empowered bots or scammers could be deployed to generate and automate sophisticated and targeted fraudulent and phishing content online, either that being an email, a professional profile, or deceptive Terms and Conditions for fake e-commerce websites. Hence, relying on outdated datasets may limit the effectiveness of detection models when applied to current or evolving threats. At the same time, the static nature of previously used datasets does not capture the latest trends and variations of various online fraud techniques and activities.
To address these limitations, recent studies have started leveraging more dynamic and real-time data sources. Regarding automated phishing detection, recent studies used user-reported phishing URLs from social networks, as well as data from Telecom and Security organizations. For instance, studies published in 2023 utilized data from sources like Twitter, Facebook, and specialized security institutions bitaab2023beyond ; saha2023phishing ; saha2023phishing ; bitaab2023beyond ; nakano2023canary ; janet2022real ; mehndiratta2023malicious ; yu2024efficient ; liang2019using .
At the same time, recent works on automated phishing email detection have utilised user-reported emails, providing a real-time perspective on phishing threats gallo2019identifying . Genc and Jianggenc2021understanding ; jiang2024detecting used emails from their personal or professional spam folders, capturing a more realistic and up-to-date snapshot of phishing attacks. Mehdi et al.mehdi2023adversarial took an innovative approach by incorporating various techniques to develop their dataset; GPT-2 generated synthetic phishing emails, along with tools like TextAttack, TextFooler, and PWWS. This approach not only provides a diverse dataset but also ensures that the model is robust against sophisticated phishing techniques. Janez et al. janez2021trustworthiness used data from the SPAM Archive [^28] http://untroubled.org/spam/, a continuously updated repository of spam emails.
Similarly, several studies on automated phishing SMS detection have extended their datasets by combining pre-existing publicly available datasets with additional sources to improve the robustness and generalizability of their models, like incorporating additional data from Fake Base Stations zhang2020lies , emails and YouTube comments vinothkumar2022detection , user reported data for research lai2022semorph or on Twitter tang2022clues . Notably, Timko et al. timko2023commercial proposed a platform, SmishTank, where users can post phishing SMS messages, creating an ongoing and up-to-date repository for researchers. We also observed studies using data from specialized security agencies seo2024device and mobile security services liu2021detecting , which offer a more targeted collection of smishing examples.
On Fake Review detection, one study bevendorff2024product used YouTube transcripts to interpret false review exaggeration, showcasing an innovative approach to identifying fraudulent content in multimedia contexts. Others have adopted similar techniques and developedtheir own data collection methodologies to collect reviews from sources like App stores, e-commerce websites, and location and travel research platforms.
The overwhelming amount of studies focus on Fraudulent Recruitment detection using the same dataset. In contrast, three studies employed custom crawlers to gather data from various job posting websites in different countries, providing a more diverse and up-to-date perspective. Mahbub et al. mahbub2022online collected data from job posting sites in the UK, Tabassum et al. tabassum2021detecting gathered job postings from Bangladeshi sites, and Zhang et al. zhang2023orfpprediction sourced data from Chinese job sites. The use of up-to-date data collection from these sources offer some advantages. For one, by scraping data from active job posting sites, these studies can capture the most recent and relevant data, reflecting current fraudulent practices. Secondly, collecting data from multiple sources across different regions provides a richer and more varied dataset, which can enhance the robustness and generalizability of detection models. Finally, custom datasets often include a wider variety of job postings, including niche or less common types of employment scams, which can be critical for developing more comprehensive detection systems.
Overall, combining publicly available datasets with recent data from other sources, such as social media, user reports, and specialized agencies, may significantly enhance the robustness and relevance of detection models. This approach could ensure that models are exposed to a wider variety of fraud tactics and can adapt to new threats more effectively.
There are various advantages of using such dynamic data sources:
  • Real-time Updates: Social networks and security organizations provide data that is continuously updated. This ensures that the detection models are trained on the most recent phishing URLs, making them more robust against new and emerging threats.
  • Diverse Data: User-reported data from social networks and institutions often include a wide variety of phishing techniques and strategies. This diversity enhances the model’s ability to generalize and detect a broader range of phishing attacks.
  • Early Detection: These sources can help in the early detection of new phishing campaigns. Social networks, in particular, can act as early warning systems where new phishing URLs are often first reported.
  • Enhanced Relevance: Data from Telecom and Security organizations are often more relevant to current threats and can include targeted phishing attacks that are not present in older datasets.
However, despite the advancements in data collection methods, there are still gaps. For instance, the data sources for various works on various online fraud categories were not clearly stated within the manuscripts. This lack of transparency can hinder reproducibility and the ability to compare results across different studies.
5.2 Methodologies
The methodologies employed across various studies on phishing and fraudulent activities involve a wide range of AI that incorporate NLP, machine learning and deep learning techniques. Most of these techniques involve using extracting features using NLP techniques and then applying supervised machine learning (i.e. labelled data) and deep learning algorithms to build binary classifiers.
For phishing URLs, Machine Learning and Deep learning algorithms such as CNN, ANN, KNN, LSTM, NB, RF, DT, SVM, and XGBoost were commonly used, often with URL feature extraction through NLP methods like character counts and n-grams. We also found various works that applied hybrid approaches where they combined multiple methodologies, demonstrating strong detection capabilities. Similarly, phishing email detection heavily relied on NLP for feature extraction (e.g., LDA, BERT) followed by AI models like RF, NB, and SVM for classification. Similar techniques were applied for phishing SMS detection.
In vishing (phishing phone calls), transcript analysis primarily conducted using NLP and AI models, with some studies examining deepfake voice detection. User reports on phishing activities utilized models like BERT and RF, while fake reviews often involved sentiment analysis with methods like VADER.
Fraudulent job postings detection involved both use of machine learning and deeep learning models such as LR and Bi-LSTM. Romance fraud detection used sentiment detection methods and machine learning models (e.g. RF) and deep learning models (e.g. LSTM). For fraudulent investment and crypto manipulation, classic machine learning models like DT, XGBoost, and SVM were employed. Studies on fraudulent e-commerce and crowdfunding leveraged advanced NLP and machine learning techniques, including GPT-4 and LR, respectively.
Although the research and detection methodologies applied by the detected literature perform well, they are not without theirlimitations. Overall, popular machine learning algorithms like RF and SVM often rely heavily on the quality of features extracted, which can be labor-intensive and may miss out on subtle indicators when dealing with large set of data.
In addition, complex models based on deep learning techniques like Bi-LSTM with VGG or hybrid approaches can be computationally expensive and difficult to implement in real-time systems due to the large amount of data processing resources required.
Notably, models developed that work well for one kind of fraud, might not generalize well on different fraud activities. For example, SMS messages are typically short, providing limited data for accurate feature extraction and classification. In addition, natural language processing techniques used may struggle to capture the features for extracting the understanding the context (semantics) or syntax of phishing content or , leading to potential false positives/negatives (i.e. incorrect misclassificaitons).
Models trained on specific languages or datasets may not perform well on emails in other languages or different styles. Many of the challenges that may arise regarding the trained AI models, are often result of the poor quality of data. More specifically, effective feature extraction is critical but can be difficult due to the varied nature of how natural language is used. At the same time, the textual content used in online fraud activities - such as fraudulent email, sms, or job posting - keeps evolving, making it challenging for static models to remain effective over time.
Although collecting data from various websites, forums, social networks, and telecommunication operators for online fraud detection is invaluable, at the same time, different platform hold data inconsistently, or have unique features and user behaviors, complicating model generalization. Also, identifying fraudulent activities in real-time is challenging due to the dynamic nature of these scams and the lack of real-time observance in the various data sources.
Overall, while these methodologies offer powerful tools for detecting phishing and fraudulent activities, they have challenges related to feature extraction, model complexity, generalizability, and data quality. Finally, one of the major issues higlighted in many of these studies is that these models are using supervised machine learning models that require labeled data. Creating labeled data is often challenging and time-consuming. As a result, researchers have been using existing labeled data, which may become less effective as fraud is evolving over time (e.g. content of phishing emails changing).
5.3 Recommendations
Datasets. The reliance on older, established datasets for training AI-based models is a double-edged sword. While they offer a solid foundation for model development and are used for omparative analysis, their static nature may limit their effectiveness in detecting evolving or emerging fraud trends, hence limiting their effectiveness for detecting new types of fraud. Therefore, there is a strong case for incorporating more dynamic and diverse data sources. Recent studies that use custom crawlers to gather data from a variety of online platforms that focus on on a range of fraud types exemplify best practices in this area. These approaches provide real-time, relevant data that can significantly improve the adaptability and accuracy of detection models. Going forward, it is recommended that researchers consider combining established datasets with freshly collected data to create more robust and resilient AI models.
Methodologies Used. Overall the majority of works used stand-alone machine learning and deep learning models for the detection of online fraud. In many cases, NLP use for feature extraction was under-reported or ignored. Work on online fraud activities that involve textual data, should utilize more sophisticated NLP techniques such as transformer-based models (e.g., BERT, GPT) for deeper semantic understanding and better context handling. Although LLMs on their own have their limitations such as generating hallucinated or inconsistent results, they are extremely powerful for context extraction.
Recent works demonstrate hybrid models, where they combine different AI and NLP methodologies to leverage their strengths, which is performing very well. Most existing studies use supervised machine learning models that require labeled data to detect fraudulent activities. Due to challenges in obtaining new labeled data, researchers often rely on existing datasets which may not capture the content of new techniques and tactics employed by scammers. To address these challenges, further exploration of active learning, semi-supervised learning and anomaly-based models that rely on small amount of labelled data or no labelled data is needed. For example, unsupervised or semi-supervised anomaly detection techniques could be studied to identify outliers and novel fraud patterns that may not be present in the training data. Finally, we observed that almost non of these models had real-time applications. There should be a shift in focus where works attempt to optimize models for real-time processing to ensure timely detection and mitigation of fraudulent activities.
Model Performance Reporting. While going through the detected literature for this SLR, we observed many works that were only reporting a subset of performance metrics of their models and frequently relying on accuracy. However, using accuracy on its own especially when the dataset is unbalanced can be misleading. In a dataset where one class has more observations than another (for example having fewer phishing emails compared to not-phishing emails), a model could achieve very high accuracy simply by predicting the majority class (i.e. not-phishing emails) without doing a good job on detecting the phishing emails.
Overall, it is essential that the researchers report a more comprehensive range of performance metrics beyond accuracy alone. This should include precision, recall, F1-score, AUC score or ROC curve. These metrics provide a more complete and nuanced picture of a model’s performance, especially in dealing with imbalanced datasets. In addition, there is a need to conduct and report detailed error analysis to identify common failure cases and the reasons behind them. This can help in understanding the limitations of the model and areas for improvement. Finally, models need to be cross-validated to ensure the robustness of the reported performance metrics. For example, reporting results from multiple folds of data samples can provide a more reliable estimate of the model’s performance.
Reproducibility. Many of the studies failed to clearly explain key and critical aspects of their model development. These include the features engineered and selected, methods used for extracting features, the data used, size of the dataset, partition of the data into training-test sets, and the hyper-parameters used for tuning and training the models.
To this end, we recommend researchers provide access to the code, datasets, and pre-trained models used in their studies through platforms like GitHub, GitLab, or institutional repositories,. This would help improve the reproducibility of their work. Researchers should also ensure that the methodology section is sufficiently detailed to allow others to replicate their study and model. This should include a clear description of pre-processing steps, feature engineering and selection, model hyperparameters, the training-testing data split and training protocols. Furthermore, the use of standardized frameworks and libraries for model implementation (e.g., TensorFlow, PyTorch) could improve reproducibility. Providing comprehensive documentation and setup instructions will further help others understand and reproduce the work easily.
Usability. Most of the studied papers are proof-of-concept, and as a result, the usability of AI-based models has not been addressed. The effective application and use of AI-based approaches depend on successful usability studies that enable users to develop these models into toolkits and provide user feedback. Usability goals are generally determined by efficiency, effectiveness, engagement, error tolerance and ease of use. It is thus imperative to ensure collaboration between the developers of AI based tools and practitioners. However, while the field of technology usability assessment in front line policing is growing 1894801 ; ZAHABI2018161 , there is a lack of usability studies considering the use of AI in preventive policing, including cybercrimes like online fraud.
Limitation/Bias discussion. The majority of studies do not discuss the limitations of their models or data. Researchers should clearly identify and report the limitations of the study, including any assumptions made, potential biases in the data, and limitations of the methodologies used. The overuse of existing labeled datasets could impact the performance of the models, potentially leading to issues such as over-fitting. However, issues like this are not discussed in the majority of works in our SLR. Researchers should be examining the distribution of different classes and any potential sources of bias in the data used for training and testing. Lastly, there is limited discussion on the generalizability of the models across different datasets, contexts, and evolving fraud patterns. Hence, research studies should test their models on a variety of experiments to evaluate generalization performance of the models.
6 Conclusion
In this systematic literature review, we have examined a wide range of studies focusing on the detection of various fraudulent activities using AI based models that use techniques from Natural Language Processing including machine learning and deep learning. Our goal was to examine the current state-of-the-art AI-based models used for development and training, investigate the sources of data, and assess how these models are evaluated for effectiveness in analysing and detecting fraudulent activities. Due to limited resources, we have restricted the data collection for the SLR to the years 2019-2024. The studies we identified highlight a focus on a wide range of fraudulent activities, with particular attention given to phishing attacks. However, there is growing interest towards the use of more advanced AI like ChatGPT for creating deceptive content as well as a tool that can be used for scam-baiting.
While significant attention have been given to building classification models that could be used to detect fraudulent activities, particularly with hybrid and advanced NLP techniques and deep learning, including LLMs , there remains a considerable room for improvement. The key AI-model development areas that require attention include performance reporting, reproducibility and transparency. Providing detailed performance reporting will help compare and evaluate different models. Improving reproducibility requires ensuring that the studies can be replicated and there is sufficient content for others to achieve this. Increasing transparency means providing clear information on how the AI-based models work and make decisions. This will help fraud practitioners to interpret and understand the models, and mitigate any biases in AI-based models.
Furthermore, most existing models rely heavily on labelled data and supervised machine-learning techniques. Future studies should give some attention to the application of unsupervised and semi-supervised machine learning for detecting fraud. Similarly, the data sources used for training these models are not suitable for capturing dynamic nature of fraud. Future studies should investigate building AI-based models that are able to capture emerging fraudulent patterns.
Addressing these gaps is crucial for creating more robust, reliable, fair, and ethical AI-based systems to detect fraud. By considering the recommended practices, the research community can better understand, detect, and mitigate fraudulent activities and reduce victimisation, ultimately contributing to a safer and more secure digital environment.
Appendix
Table 3: Data sources and Detection Methods used for Phishing URL detection. A single asterisk (*) indicates that the data is not publicly available. UNK indicates Unclear details. Empty cells indicate missing values. P: Precision, R: Recall, A: Accuracy, F1: F1 Score, AUC: Area under the Curve.
# URL Source Collection Method Models Used Best Model Performance
P R A F1 AUC
alswailem2019detecting (2019) phishtank.org* Custom crawler RF RF 0.98
rao2019detection (2019) phishtank.org and Alexa* UNK J48, RF, SMO, LR, MLP, BN, SVM, AdaBoost RF 0.99 0.99
almseidin2019phishing (2019) phishtank.org and openphish.com* Previous work chiew2019new BNET, NB, J48, LR, RF, MLP RF 0.98
chiew2019new (2019) Alexa, phishtank.org, openphish.com, and commoncrawl.org* UNK RF, SVM, NB, C4.5, JRip, PART RF 0.94
li2019stacking (2019) Alexa and phishtank.org* UNK SVM, KNN, DT, RF, GBDT, XGBoost, LGB, Hybrid (GBoost, XGBoost, and LightGBM) Hybrid (GBoost, XGBoost, and LightGBM) 0.97
yadollahi2019adaptive (2019) UNK UNK C4.5, AdaBoost, KNN, RF, SMO, NB Hybrid (XCS/UNK) 0.98 0.98 0.98 0.99
sahingoz2019machine (2019) phishtank.org, Yandex Search API, and GitHub Open dataset github-url-phishing and custom crawler DT, AdaBoost, Kstar, KNN, RF, SMO, NB DT 0.96 0.97 0.97
liang2019using (2019) NetLab360 and Alexa* UNK LR, SVM, LSTM LSTM 0.98 0.98
zamir2020phishing (2020) Kaggle Open dataset kaggle-url-data-kumar NB, KNN, SVM, RF, Bagging, NN Hybrid (NN, RF, and Bagging) 0.95 0.98 0.97 0.96
somesha2020efficient (2020) Alexa and phishtank.org* Previous work rao2019detection DNN, LSTM, CNN LSTM 0.99
tharani2020understanding (2020) Alexa, phishtank.org, Mendeley, openphish.com, and commoncrawl.org* Open dataset mendeley-url-data and custom crawler NB, SVC, KNN SVC, KNN
do2020malicious (2020) phishtank.org, URLHaus, Majestic, Kaggle Open datasets kaggle-url-data-anthony ; majestic-data ; urlhaus-data and custom crawler SVM, RF RF 0.98 0.97 0.99
kumar2020phishing (2020) Refer to open dataset GitHub open dataset githubURLdataset NB NB 1 0.95 0.97
saha2020phishing (2020) Kaggle* UNK open dataset MLP MLP 0.93
al2020convolutional (2020) UNK* Previous works mohammad2014predicting ; mohammad2014intelligent RF, RNN, CNN CNN 0.94 0.91
priya2020gravitational (2020) UCI ML Repository Open dataset UCI-ml-phishing-url RF, DT, ANN, KNN RF 0.95
rao2020two (2020) phishtank.org and Google Search* UNK SVM, DT, LR, RF, XGBoost, AdaBoost, ET Hybrid (RF, XGBoost and ET) 0.98
chen2020intelligent (2020) phishtank.org* UNK KNN KNN 0.98
raja2021lexical (2021) Kaggle and Canadian Institute of Cybersecurity* UNK SVC, LR, KNN, NB, RF KNN 0.98 0.98 0.98 0.98
chen2021ai (2021) scammer.info and urlscan.io* Custom crawler LGBM LGBM 1 0.96 0.98 0.97 0.98
geyik2021detection (2021) UNK* Previous work rao2020catchphish RF, DT, NB, LR RF 0.83
el2021malweb (2021) UNK* UNK LR, DT, NB LR, DT 1 1 1 1
ou2021no (2021) Alexa and cryptoscamdb.org* Custom crawler NB, SVM, KNN, RF RF 0.98 0.95 0.97 0.96
salloum2021phishing (2021) Alexa, phishtank.org, and Mendeley* Custom crawler and open dataset mendeley-url-phishing XBoost, RF, SVM, KNN, ANN, LR, DT, NBB ANN 0.96 0.97 0.97 0.97
barraclough2021intelligent (2021) phishtank.org, relbanks.com, and millersmiles.co.uk* Custom crawler ANFIS, NB, PART, J48, JRip PART 0.98 0.99 0.99 0.99 0.98
chen2022development (2022) Google Rankings and whoscall.com Custom crawler RF, DNN RF 1 0.99 0.99
li2022mui (2022) Alexa and phishtank.org* Custom crawler Bi-LSTM, Hybrid (Bi-LSTM and CNN), Hybrid (Bi-LSTM and VGG) Hybrid (Bi-LSTM and VGG) 0.96 0.96 0.96
shalke2022social (2022) who.is* Custom crawler BPNN, RBFN, SVM, NB, DT, RF, KNN NB 0.96
alkawaz2022identification (2022) UNK* UNK DT, RF RF 0.8
navyah2022ensemble (2022) Alexa, UCI, phishtank.org and Kaggle* UNK open dataset KNN, RF, DT, CBoost, LGBM, ABoost, VC CBoost 0.98 0.98
jaber2022improving (2022) phishtank.org and UCI ML Repository Custom Crawler and open dataset UCI-ml-phishing-url Hybrid (CNN) Hybrid (CNN) 0.97
vecile2022malicious (2022) Canadian Institute for Cybersecurity Open dataset cicURLdataset LSTM LSTM 0.99 0.99 0.99
gu2022ensemble (2022) Kaggle* UNK RF, KNN, XGBoost XGBoost 0.96 0.96
mandadi2022phishing (2022) pishitank.org* Custom crawler RF,DT RF 0.87
villanueva2022application (2022) GitHub Open dataset github-url-phishing LR, NB, LSTM, GRU LSTM or GRU 0.95
mohammed2022accuracy (2022) UCI ML Repository Open datasets UCI-ml-phishing-url and UNK AdaBoost, CART, GBoost, MLP, SVM, RF, NB, SEM SEM 0.98
marimuthu2022intelligent (2022) phishtank.org and Alexa* Custom crawler DT, RF RF 0.87
fernandez2022early (2022) Farsight SIE fss , spamhaus.org, and surbl.org* UNK J48, RF RF
shaiba2022hunger (2022) UNK* Previous work rao2020catchphish Hybrid (CNN and LSTM) (CNN and LSTM) 0.98 0.99 0.99 0.99 0.99
ariyadasa2022phishrepo (2022) Mendeley and previous works Mendeley open dataset mendeley-phishrepo-data and previous works lin2021phishpedia ; feng2020web2vec Hybrid DLM, Stack model, URLNet Hybrid DLM 0.93 0.93
orunsolu2022predictive (2022) phishtank.org and Alexa* UNK SVM, NB Hybrid (UNK) 0.99 0.99 0.99 0.99
pradeepa2022lightweight (2022) Canadian Institute of Cybersecurity, phishtank.org, and Kaggle* UNK SVM, RF RF 0.99 0.99 0.99
janet2022real (2022) Twitch* Twitch API XGBoost, RF, NB RF 0.93 0.93 0.93
vo2023shark (2023) phishtank.org and openphish.com Custom crawler CNN SharkEyes (CNN, W2V, GRU, Bi-LSTM) 0.94 0.94 0.95 0.94
nakano2023canary (2023) Tweets, spamhunter.io, and tweetfeed.live* Twitter APi and custom crawler Hybrid (BERT and RF) Hybrid (BERT and RF) 0.96 0.95 0.95
saha2023phishing (2023) Twitter and Meta’s crowdtangle.com* Twitter API and custom crawler UNK Pre-trained model li2019stacking 0.96 0.97 0.97 0.96
jha2023machine (2023) Kaggle* Data no longer available RF, LR, KNN RF 0.97 0.99 0.97
bitaab2023beyond (2023) reddit.com/r/Scams/ and Paolo Alto Networks* Custom crawler RF, XGBoost, SVM, FFNN BeyondPhish (RF and XGBoost and SVM and FFNN) 0.98
mehndiratta2023malicious (2023) Kaggle and Canadian Institute of Cybersecurity Open datasets kaggleURLdataset ; cicURLdataset KNN, LR KNN 0.9
jain2023support (2023) Kaggle* UNK open dataset DT, KNN, RF, SVM SVM 0.99 0.96 0.98 0.97
zin2023machine (2023) Mendeley Open dataset mendeley-url-data RF, J48, NB, KNN, LR RF 0.97 0.9 0.94
kumar2023hybrid (2023) Kaggle Open dataset kaggle-url-data DT, KNN, RF, GBoost UNK hybrid 0.98
aslam2023phish (2023) Mendeley Open dataset mendeley-url-data MLP, RF, RT, KNN, SVM RF 0.98 0.98 0.98
pathak2023classification (2023) UNK* Previous work pathak2023classification DT, KNN, SVM, NB, LR, XBoost, Aboost Hybrid (DT, SVM, LR) 0.99 0.98 0.99 0.99
jishnu2023enhanced (2023) phishTank, Kaggle, and Majestic* UNK BERT BERT 0.97 0.96 0.97 0.97
rafsanjani2023qsecr (2023) URLHaus and phishtank.org Custom crawler Custom rule based Custom rule based 0.93 0.93 0.93 0.93
kalabarige2023boosting (2023) UCI ML Reposiroty and Mendeley Open dataset mendeley-url-data ; UCI-ml-phishing-url ; mendeley-url-phishing LGBM, XGBoost, AdaBoost, CatBoost, GB, Hybrid (BMLSELM) Hybrid (BMLSELM) 0.97 0.97 0.97 0.97
ashwitha2023perception (2023) Kaggle* UNK LR, NB, DT, SVM, RF, KNN KNN 0.99
dr2023malicious (2023) Kaggle* UNK RF, XGBoost, LightGBBM RF 0.99 0.94 0.96 0.96
kundra2023identification (2023) UNK* UNK RF, AdaBoost, XGBoost, GBoost, KNN RF 0.91
jha2023intelligent (2023) phishtank.org* UNK LR, RF RF 0.93 0.79 0.96 0.85
nagy2023phishing (2023) PubMed Open dataset singh-url-phishing RF, NB, LSTM, CNN UNK
adebowale2023intelligent (2023) phishtank.org and who.is* Custom crawler LSTM, CNN, Hybrid (LSTM and CNN) Hybrid (LSTM and CNN) 0.93
yu2024efficient (2024) Zhejiang Mobile Innovation Research Institute* UNK MBERT, XGBoost, LBoost, LSTM, NB, LR, RF, SVM, KNN MBERT 0.94 0.94 094
Table 4: Data sources and Detection Methods used for Phishing Email detection. A single asterisk (*) indicates that the data is not publicly available. UNK indicates Unclear details. Empty cells indicate missing values. P: Precision, R: Recall, A: Accuracy, F1: F1 Score, AUC: Area under the Curve.
# URL Source Collection Method Models Used Best Model Performance
P R A F1 AUC
salihovic2019role (2019) UNK* UNK RF, KNN, ANN, SVM, LR, NB RF 0.97
gallo2019identifying (2019) Spam emails received by a company* UNK GNB, DT, SVM, NN, RF RF, SVM 0.92 0.97 0.89
markova2019classification (2019) cs.cmu.edu Open dataset enron-email-dataset RF, KNN, SVM, DT RF 0.92 0.94 0.91
al2020email (2020) aclweb.org and previous work Open dataset acl-email-dataset and previous work almeida2011contributions NB, Dt, RF, SVM SVM 0.98 0.97 0.98 0.97
rahmad2020performance (2020) Spam emails received by a company* UNK Clustering Clustering 0.89
islam2021spam (2021) Open datasets* UNK LR, SVM, RF, XGBoost XGBoost
bhatti2021email (2021) cs.cmu.edu Open dataset enron-email-dataset LSTM LSTM 0.97
stojnic2021phishing (2021) UNK UNK Various topic modelling N/A N/A N/A N/A N/A N/A
genc2021understanding (2021) Author’s spam folder* Custom LDA, Jaccard N/A N/A N/A N/A N/A N/A
jonker2021using (2021) UNK* UNK RNN, LSTM, CNN, BERT UNK
janez2021trustworthiness (2021) Questionnaires and untroubled.org/spam Open dataset [^29] https://untroubled.org/spam/ NB, SVM, RF, LR NB 0.88 0.8
venugopal2022detection (2022) UNK* Custom crawler BOW (Rule based) BOW (Rule based) 0.99 %
singh2022spam (2022) Kaggle Open dataset kagge-email-dataset CBoost, LR, DT, RF, GNB, SVM, KNN, XGBoost, LGBM, AdaBoost CBoost 0.97 0.96 0.97
livara2022empirical (2022) Kaggle Open dataset kagge-email-dataset2 RF, NB, SVM, AdaBoost, LR RF 0.99
al2022digital (2022) Previous work* Previous work hina2021sefaced RF, LR, SVM, MNB RF, LR, SVM, 0.95 0.95 0.95
saka2022context (2022) GitHub, monkey.org, cs.cmu.edu Open datasets github-email-dataset ; jose-monkey-email-dataset ; enron-email-dataset K-Means, DBSCAN, and Agglomerative Clustering Agglomerative Clustering N/A N/A N/A N/A
mughaid2022intelligent (2022) UNK UNK SVM, DT, LR, DNN, RF DT 1 1 1 1
ismail2022efficient (2022) UCI ML Reposiroty* UNK NB, SVM, KNN, J48, DT DT 0.98
janez2023review (2023) Previous work, cs.cmu.edu, spamassassin.apache.org, and csmining.org [^30] Data link broken Previous work androutsopoulos2000learning ; cormack2007spam , open datasets enron-email-dataset ; spamassasin-email-dataset , and UNK NB, SVM UNK
mehdi2023adversarial (2023) Synthetic data Data generated using various techniques github-generated-email-dataset ALBERT, RoBERTa, BERT, DBERT, SQ, YOSO ALBERT 0.94 0.95
ramprasath2023identification (2023) Kaggle* UNK RNN, LSTM, CNN RNN 0.99 0.92 0.99 0.95
kushwaha2023analysis (2023) UCI ML Repository Open datasetuci-sms-dataset BERT BERT 0.95 0.93 0.98 0.94
saini2023machine (2023) UCI ML Repository* Previous work saini2023machine SVM, RF, NB RF 0.95
jena2023malicious (2023) Previous work* Previous work yerima2022semi KNN, NB, DT, RF, SVM, LR, XGBoost, BERT BERT 0.97 0.97 0.97 0.97
mittal2023blockage (2023) UCI ML Repository* UNK CatBoost CatBoost 0.97 0.96 0.96 0.97 0.99
bera2023towards (2023) Previous work, Kaggle, and monkey.org Previous work el2020depth ; sakkis2003memory ; metsis2006spam and open datasets jose-monkey-email-dataset ; kaggle-email-dataset5 Various topic modelling N/A N/A N/A N/A N/A N/A
emmanuel2023information (2023) Kaggle Open dataset kagge-email-dataset3 MLP, DT, LR, RF, KNN, SVM MLP, SVM 0.99 0.99 0.99 0.99
chataut2024can (2024) Kaggle Open dataset kaggle-email-dataset4 GPT-3.5, GPT-4, Custom (CyberGPT) Custom (CyberGPT) 0.97
jiang2024detecting (2024) UNK* UNK GPT-3.5, GPT-4 UNK
Table 5: Data sources and Detection Methods used for Phishing SMS detection. A single asterisk (*) indicates that the data is not publicly available. UNK indicates Unclear details. Empty cells indicate missing values. P: Precision, R: Recall, A: Accuracy, F1: F1 Score, AUC: Area under the Curve.
# URL Source Collection Method Models Used Best Model Performance
P R A F1 AUC
jain2019feature (2019) Previous work* Previous work almeida2011contributions SVM, LR, NN, NB, RF RF 0.98
zhang2020lies (2020) 360 Mobile Safe* UNK SVM, NB, LR, RF SVM 0.96 0.96 0.96
liu2021detecting (2021) 360 Mobile Safe* UNK LR, DT, NB, SVM LR 0.93 0.93 0.93
ulfath2021hybrid (2021) UCI ML repository Open dataset uci-sms-dataset CNN, GRU, MLP, SVM, XGBoost, Hybrid (CNN, GRU) Hybrid (CNN, GRU) 0.99 0.96 0.98
lai2022semorph (2022) https://www.datafountain.cn/* Custom crawler CNN, BERT, RoBERTa, ChineseBERT Hybrid (Semorph/UNK) 0.96 0.84 0.89
tang2022clues (2022) Twitter* Twitter API Custom/UNK Custom/UNK 0.98 0.97 0.97
jain2022sms (2022) UCI ML repository* UNK SVM, NB, LR, DT SVM 0.96 0.93 0.98 0.95
vinothkumar2022detection (2022) Kaggle and YouTube* UNK NB, DT, KNN NB 0.97
abid2022spam (2022) Kaggle Open dataset kaggle-sms-dataset LR, SVC, RF, NB, GBM RF 0.99 0.95 0.99 0.97
timko2023commercial (2023) smishtank.com* Custom crawler Various NLP methods N/A
addanki2023safeguarding (2023) Kaggle and inaccessible website Open dataset kaggle-sms-dataset and custom crawler LinearDA, QDA, SVM, PCA, NB SVM 0.97
dharani2023spam (2023) Kaggle Open dataset kaggle-sms-dataset KNN, NB, RF, SVC, ETC, LR, XGBoost, AdaBoost, GBDT, DT, NBB 1 0.95
zhang2023bert (2023) Previous work* Previous work zhang2020lies BERT-GCN BERT-GCN 0.92 0.96 0.93
al2023multi (2023) UCI ML repository* UNK LSTM, CNN, RF, Hybrid (various), BERT, LSTM, XGBoost Hybrid (CNN, LSTM) 0.99 0.99 0.99 0.99
gandhi2023sms (2023) Kaggle Open dataset kaggle-sms-dataset DNN, LSTM DNN 0.95
siagian2023improving (2023) UCI ML repository Open dataset uci-sms-dataset LSTM, GRU, NB, BERT BERT 0.99 0.99
kohilan2023machine (2023) Kaggle and Mendeley* UNK SNN, RNN, CNN CNN 0.99
mishra2023dsmishsms (2023) UNK UNK BPA, RF, NB, DT BPA 0.97 0.98
agrawal2023effective (2023) UNK UNK NB, RF, ETC ETC 0.99 0.96
seo2024device (2024) Korean Internet and Security Agency UNK NB, RF, LGBM, CNN, KoELECTRA CharCNN 0.99 0.99
Table 6: Data sources and Detection Methods used for Fake User detection. A single asterisk (*) indicates that the data is not publicly available. UNK indicates Unclear details. Empty cells indicate missing values. P: Precision, R: Recall, A: Accuracy, F1: F1 Score, AUC: Area under the Curve.
# Data Collection Method Models Used Best Model Performance
P R A F1 AUC
ali2019detect (2019) Twitter user profiles Open dataset yang2013empirical GCNN, MLP, BP GCNN 0.94
albayati2019identifying (2019) Facebook user profiles* UNK ID3, KNN, SVM ID3 0.98 0.98 0.97
yue2019madafe (2019) Twitter User Profiles Open dataset WU2018265 SVM, RF, MADAFE (NN and LR) MADAFE
venkatesan2019graph (2019) Twitter and Facebook (UNK)* UNK HDBSCAN HDBSCAN N/A N/A N/A N/A N/A
raj2020multi (2020) Tweets* Twitter API KNN, RF, NB, DT RF 0.95
zhang2021social (2021) Sina Weibo User profiles* Custom crawler CatBoost, RF CatBoost 0.87
bebensee2021leveraging (2021) Twitter user profiles Open dataset Cresci2018 NB, QDA, SVM, KNN, RF, NN RF 0.87 0.88 0.94
anklesaria2021survey (2021) Instagram user profiles* Instagram API RF, AdaBoost, MLP, ANN, SGD RF 0.99 0.98 0.98 0.98 0.98
shreya2022identification (2022) Facebook user profiles* Manual collection ANN, SVM, RF ANN 0.96
das2022effecient (2022) Instagram user profiles* UNK LR, KNN, SVM, RF, NB RF 0.99 0.97 0.94 0.98
rovito2022evolutionary (2022) Twitter user profiles Open dataset feng2021twibot GA, GP GP 0.76 0.78
shukla2022tweezbot (2022) Twitter user profiles Open dataset jain_kaggle SVM, CNB, BNB, MP, DT, RF TweezBot (Unclear) 0.99 0.93 0.98 0.99
na2023evolving (2023) YouTube video and user metadata and comments* YouTube API Sentence-BERT, RoBERTa, YouTuBERT YouTuBERT (LLM and DBSCAN) 0.63 0.81 0.90 0.71
haq2023spammy (2023) List of names* UNK NB, KNN, SVM, LR, RF NB 0.94 0.94 0.95 0.94
gangan2023detection (2023) Twitter user profiles* UNK NB, DT, NN, Ensemstack Ensemstack 0.98
fathima2023ann (2023) Instagram user profiles* UNK ANN ANN 0.74
nikhitha2023fake (2023) Instagram user profiles* UNK LR, DT, RF RF 0.9
singh2023safe (2023) Twitter user profiles and tweets* Twitter API LR LR 0.93 0.93 0.93 0.93
Table 7: Data sources and Detection Methods used for Fraudulent Recruitment detection. The asterisk (*) indicates that the data is not publicly available. Empty cells indicate missing values. P: Precision, R: Recall, A: Accuracy, F1: F1-Score, AUC: Area Under the Curve.
# Job postings source Collection Method Models Used Best Model Performance
P R A F1 AUC
lal2019orfdetector (2019) Kaggle Open dataset emscad-dataset J48, LR, RF Ensemble 0.95 0.94
ranparia2020fake (2020) Kaggle Open dataset emscad-dataset GLoVE GLoVE 0.99
nasser2021online (2021) Kaggle Open dataset emscad-dataset ANN ANN 0.91 0.96 0.93
vo2021dealing (2021) Kaggle Open dataset emscad-dataset LR LR 0.89 0.92 0.96
li2021exploratory (2021) Kaggle Open dataset emscad-dataset LightGBM, LR, DT, XGBoost, AdaBoost LightGBM 0.93 0.94 0.95 0.93
tabassum2021detecting (2021) job.com.bd, bdjobstoday, and deshijob* Custom crawler LR, AdaBoost, DT, RF, VC, LGBM, GB LightGBM or GBoost 0.95
habiba2021comparative (2021) Kaggle Open dataset emscad-dataset KNN, RF, DT, SVM, NB, DNN DNN 0.97
amaar2022detection (2022) Kaggle Open dataset emscad-dataset RF, LR, SVM, ETC, KNN, MP ETC 0.99
bhatia2022detection (2022) Kaggle Open dataset emscad-dataset KNN, RF KNN 0.79 0.73 0.98 0.76
mahbub2022online (2022) SEEK, Glassdoor, Indeed, and Gumtree job postings* Custom crawler RF, JRip, NB, J48 RF 0.82 0.69 0.91
nessa2022recruitment (2022) Kaggle Open dataset emscad-dataset GRU GRU 0.93
prathaban2022verification (2022) Kaggle Open dataset emscad-dataset LR, NB, MLP, KNN, RF, DT, Adaboost, GB, NLP RF 0.98 0.97 0.97 0.98
pandey2022effective (2022) Kaggle Open dataset emscad-dataset RF, SVM, Bi-LSTM Bi-LSTM 0.98 0.98
yang2023improved (2023) Kaggle Open dataset emscad-dataset SVM, NB, RF, Bi-LSTM, LR RF
reddy2023web (2023) Kaggle Open dataset emscad-dataset RF, XBoost, LightGBM, CatBoost, DT XGBoost 0.95 0.9 0.96 0.92
santhiya2023fake (2023) Kaggle Open dataset emscad-dataset RF, SVM, NB, Ensemble RF 0.98
sofy2023intelligent (2023) Kaggle Open dataset emscad-dataset RF, NB, SVM, DT, KNN RF 0.97
nanath2023investigation (2023) Kaggle Open dataset emscad-dataset LR, DT, RF, NB, GLM GLM 0.96 0.78 0.86 0.98
ullah2023smart (2023) Kaggle Open dataset emscad-dataset AdaBoost, XGBoost, RF, Voting AdaBoost 0.99 0.97 0.98 0.98
zhang2023orfpprediction (2023) Boss Zhipin, Liepin, 51job* Custom crawler NB, XGBoost, SVM, LightGBM, DT, RF DRLM (DT and RF and LightGBM) 0.98 0.94 0.92
Table 8: Data sources and Detection Methods used for Fake Review detection. A single asterisk (*) indicates that the data is not publicly available. UNK indicates Unclear details. Empty cells indicate missing values. P: Precision, R: Recall, A: Accuracy, F1: F1-Score, AUC: Area Under the Curve.
# Data Collection Method Models Used Best Model Performance
P R A F1 AUC
gupta2020detecting (2020) Amazon reviews* Custom crawler SVM, LR, RF, DT, GNBSGD, KNN, 3LP, 4LP, XGBoost 3LP 0.98 0.98 0.98 0.98
furia2020tool (2020) Amazon reviews* Custom crawler SVM, KNN, NB, Ensemble Ensemble 0.81 0.81 0.81 0.81
wang2021modeling (2021) Yelp and JD.com reviews Open datasetyelpzip ; yelpnyc and JD.com custom crawler GraphSAGE, Cluster-GCN, HGT, C-FATH (Custom) C-FATH (Unclear)* 0.68-0.87 0.95-0.97
chandana2021analyzing (2021) Amazon reviews Open dataset liu2019method RF RF* 1 0.85 0.98
javed2021fake (2021) Yelp reviews* UNK CNN, SVM, LR, MLP CNN 0.93 0.92 0.92
balakrishna2022identifying (2022) Yelp reviews Open dataset asghar2016yelp ; ott2013negative WaveNet, LDA WaveNet, LDA N/A N/A N/A N/A N/A
deekshan2022detection (2022) Amazon reviews* UNK BERT, VADER, LSTM, WordNet, SGD, SVM, LR LR 0.81
rangari2022empirical (2022) Amazon hotel reviews* UNK KNN, NB, SVM SVM* 0.93
tushev2022domain (2022) Smartphone App reviews* Web Scraping LDA, keyATM keyATM N/A N/A N/A N/A N/A
obie2022violation (2022) Smartphone App reviews Open dataset eler2019android ; obie2021first SVM, DT, NN, LR, GBT SVM 0.94 0.84 0.89
yugeshwaran2022rank (2022) Google Play reviews* Custom crawler DT, RF, MLP MLP 0.97
rangar2022machine (2022) Amazon reviews* UNK CNN, SVM, NB CNN 1 1 1
tufail2022effect (2022) Hotel reviews Open dataset ott2013negative ; yelp-dataset SVM, KNN, LR SKL (SVM and KNN and LR) 0.95
harris2022combining (2022) Yelp reviews Open dataset yelpzip Bi-LSTM Bi-LSTM 0.89
akshara2023small (2023) Amazon book reviews* UNK SVM, LR LR 0.86
singh2023deep (2023) Reviews Open dataset ott2013negative and UNK CNN, LSTM, KNN, NB, SVM, W2V CNN, LSTM 0.93
iqbal2023efficient (2023) Amazon reviews* Custom crawler AdaBoost, RF, Lr, SVM, KNN RF 0.99 0.99 0.99 0.99
ashraf2023fake (2023) Yelp reviews* UNK SVM, MLP, CNN, LR CNN 0.85 0.85 0.85 0.85
silpa2023detection (2023) Yelp reviews* UNK NB, LR, SVM, DT SVM 0.96 0.98 0.97
pengqi2023unmasking (2023) Yelp reviews Open dataset yelpzip GPT-3, BERT, RF, XGBoost GPT-3 0.73 0.64 0.68 0.75
ganesh2023implementation (2023) Undefined reviews* UNK ANN, CNN, LR, SVM, NB, KNN, RF, DT, SGD LR 0.89
thilagavathy2023fake (2023) Hotel reviews* UNK SVM SVM
bevendorff2024product (2024) Product reviews* YouTube API SVM,LR LR 0.74 0.99 0.85 0.95
Acknowledgements
The authors would like to thank the reviewers for their feedback and reviews.
Funding
This work was funded by the Dawes Trust.
Abbreviations
PRISMA-ScR: Preferred Reporting Items for Systematic reviews and Meta-Analyses extension for Scoping Reviews; NLP: Natural Language Processing; AI: Artificial Intelligence; ML: Machine Learning; SLR: Systematic Literature Review.
Availability of data and materials
The data extracted from the academic papers included in this review will be released with the full accepted paper.
Ethics approval and consent to participate
Not applicable.
Competing interests
The authors declare that they have no competing interests.
Consent for publication
Not applicable.
Authors’ contributions
AP and NT drafted the final manuscript. AP conducted the updated academic literature review, designed and conducted the grey literature review, extracted data from the literature, and analysed and interpreted the results of all aspects of the scoping review. SJ and ED provided substantial feedback on the review process and editing of the document. EM contributed to the coding process. AM and SL contributed to the editing process.
Authors’ information
1 Security and Crime Science, University College London, London, United Kingdom.
2 Humanities and Social Sciences, Anglia Ruskin University, London, United Kingdom.
3 Advanced Research Computing Centre, University College London, London, United Kingdom.